### v3

# thanks: (No relevant financial or non-financial competing interests. Derived from preprint [23].)

Abstract
The multi-armed bandit (MAB) problem is a widely studied model in the field of operations research for sequential decision making and reinforcement learning. This paper mainly considers the classical MAB model with the heavy-tailed reward distributions. We introduce the extended robust UCB policy, which is an extension of the pioneering UCB policies proposed by Bubeck et al. [5] and Lattimore [21]. The previous UCB policies require the knowledge of an upper bound on specific moments of reward distributions or a particular moment to exist, which can be hard to acquire or guarantee in practical scenarios. Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p$ and $q$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of the UCB policies for the heavy-tailed reward distributions.
keywords:
frequentist multi-armed bandit problems; the light-tailed reward distributions; the heavy-tailed reward distributions; upper confidence bounds
1 Introduction
The multi-armed bandit problem is an important topic in operations research where a player aims to identify and select the best arm to operate such that the long-term expected reward is maximized. The player’s action at each time involves balancing the trade-off between exploration and exploitation, i.e., choosing an arm to learn its statistics vs. choosing an arm appearing to be best based on past observations. Under resource constraints, the way we acquire knowledge becomes crucial for the efficiency in maximizing the long-term reward in modern AI systems [31]. While knowledge can be generally treated as data (usually not free in a strict sense), the struggle between exploration (sampling) and exploitation (consuming) as the prototypical dilemma in reinforcement learning is expressly reflected in an MAB model [28, 20, 3, 24, 5]. Specifically, the original stochastic model of the multi-armed bandit problem considers a bandit machine with several arms; each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or, agent). Each reward distribution has a mean, say $\theta_{i}$. The player who has limited knowledge about the underlying reward distributions (no knowledge about the means $\{\theta_{i}\}$), has to choose only one of these arms to play and accrue a reward at each time step, given the reward history so far observed. The target of the player is to maximize the rate of gaining reward as the time horizon $T$ grows. In many applications of online decision-making problems, such as Internet advertising and recommendation systems, the MAB model is widely assumed with numerous algorithms proposed for efficient reinforcement learning, e.g., Zhao [38], Lattimore and Szepesvári [22]. In a recommendation system, the vendor desires to know which item an online customer is looking for with the most interest. When we suggest a product and the customer makes a purchase, we receive a positive reward. Conversely, if the customer decides not to buy the recommended item, our reward may be either zero or negative, reflecting a potential waste of resources. Therefore the target of a recommendation system is to guess which items are most wanted by each customer, and the problem can be translated into the MAB model directly.
To address the key issue in the conflict between exploration and exploitation, we note that it is impossible for the player to exactly know the expected values of the reward distributions within any finite horizon of time. The only information the player possesses is the (local) reward history, and the player has to learn from this to make a decision at each time step. Based on the limited historical information, the player has to choose between two options. The first option is to explore the arms with a lower average reward, as there is always a possibility that the best arm may not give the best average reward in the past sample path. The second option is to exploit the arm with the highest average reward offered so far. Robbins [28] considered a simple two-armed bandit problem and showed that there is a policy to achieve the best average reward asymptotically as time goes to infinity. Many years later, Lai and Robbins [20] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the benchmark case that the player already knows which arm has the highest reward mean from the very beginning. Given a set of $K$ arms with reward distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$ with means $\{\theta_{i}\}_{i=1}^{K}$ and a policy $\pi$, the regret ~$\tilde{\mathcal{R}}_{\pi}^{\mathcal{F}}(T)$ is defined as (without loss of generality, we can assume that $\theta_{1}\geq\theta_{i}$ for $1\leq i\leq K$)
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=\sum_{\theta_{i}<\theta_{1}}E[s_{i}(T)]\Delta_{i},$ (1)
where
$\Delta_{i}=\theta_{1}-\theta_{i},$ (2)
and $s_{i}(T)$ denotes the number of times that arm $i$ was chosen during time steps 1 to $T$ under policy $\pi$. Certainly, a lower regret growth rate indicates better performance of a policy, and also a higher efficiency of learning. Furthermore, any sub-linear regret order implies the best expected time-average reward as $T\rightarrow\infty$.
Strikingly, Lai and Robbins [20] proved an asymptotic lower bound on regret, with the growth rate of $O(\log T)$, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown. They also proposed an asymptotically optimal policy for a family of distributions, under which the best regret can be achieved asymptotically, both the logarithmic order and the leading coefficient. Later Burnetas and Katehakis [6] proved a generalized result under non-parametric settings. Suppose $\mathcal{H}$ is a set of distributions on $\mathbb{R}$. We call a bandit $\mathcal{F}$ (considered as a set of reward distributions of all arms) non-trivial, if there is at least one distribution (re-index it to $1$ if necessary) $F_{1}\in\mathcal{F}$ satisfying $\theta_{1}>\theta_{i}$ for all $i\in\{2,\dots,K\}$. They proved that for any policy $\pi$ s.t. for each fixed nontrivial $\mathcal{F}\subset\mathcal{H}$ (this policy condition is called consistency over $\mathcal{H}$)
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=o(T^{a}),\ \forall a>0,$ (3)
we have that
$\lim\inf\frac{\mathbb{E}[s_{i}(T)]}{\log T}\geq\sup\left\{\frac{1}{\text{KL}(F_{i},Y)}:Y\in\mathcal{H}\ \text{and}\ \mathbb{E}Y>\theta_{i}\right\},$ (4)
where $\text{KL}(\cdot,\cdot)$ denotes the Kullback-Leibler distance between two distributions. This theorem demonstrates that for any consistent strategy under a given family of distributions, the regret will not be lower than the logarithmic growth. Auer et al. [3] proposed a class of policies named and featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $[a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player. Their UCB1 and UCB1-Tuned polices perform very well for the classical MAB problem with a theoretically proven regret bound for any finite time horizon. The UCB policies are widely applied in machine learning development, e.g., the recommendations systems [32, 16], and in various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [19], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [15].
1.1 Heavy-tailed MAB Model
The class of heavy-tailed distributions is a concept in contrast to the light-tailed case. Light-tailed (or, sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation, while heavy-tailed distributions have a relatively high probability of producing a large deviation. There are several popular light-tailed probability distributions, such as Bernoulli, Gaussian, Laplacian, and Exponential. Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions [10]. Formally, a random variable $X$ is called light-tailed if there exists some $u_{0}>0$ such that its moment-generating function is well-defined on $[-u_{0},u_{0}]$ [10]:
$M(u):=\mathbb{E}[\exp(uX)]<\infty,\ \ \forall|u|\leq u_{0}.$ (5)
The above condition is equivalent to [10]
$\mathbb{E}[\exp(u(X-\mathbb{E}X))]\leq\exp(\zeta u^{2}/2),\ \ \forall|u|\leq u_{0},\ \forall\zeta\geq\sup_{|u|\leq u_{0}}{M^{(2)}(u)},$ (6)
where $M^{(2)}(\cdot)$ denotes the second derivative of $M(\cdot)$. Note that the above upper bound on $M(u)$ has a form of the moment generating function of the Gaussian distribution, thus the name “sub-Gaussian” has been adopted.
The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution might take large values with higher probabilities, thus its moment-generating function does not exist (even locally). Compared to the light-tailed distributions, the heavy-tailed are harder to learn in terms of the rank of their means. Bubeck et al. [5] proposed the robust UCB algorithm, under the assumption that there exists a known $p>1$ such that $\mathbb{E}[|X_{i}|^{p}]$ exists and is upper bounded by a known parameter, for all $1\leq i\leq K$. Robust UCB shows significant progress for the heavy-tailed MAB model by achieving the optimal logarithmic order of regret growth. Later on, Lattimore [21] developed a new UCB policy, which is referred to as the scale free algorithm. This algorithm removes the general assumption of a known upper bound on the $p$-th moment. Instead, it allows the variance and the squared fourth moment to scale freely as long as their ratio (kurtosis) is bounded by a known parameter. But this assumption restricts $p$ to be no less than $4$ (thus cannot be too heavy-tailed). In this paper, we extend Lattimore’s work to any $p>1$ with a tighter regret upper bound in the case of $p=4$ and low discrimination where arm means are close and learning the arm rank becomes challenging.
From this point on, we assume that the player knows a constant $C_{p,q}$ for some $p$ and $q$ ($p>q>1$) such that
$\mathbb{E}[|X-\mathbb{E}X|^{p}]\leq C_{p,q}\left(\mathbb{E}[|X-\mathbb{E}X|^{q}]\right)^{p/q}.$ (7)
The above inequality is a direct generalization of the assumption on kurtosis in Lattimore [21] where $p=4$ and $q=2$. For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function [2]
$F(x)=\begin{cases}1-\left(1+\left(\frac{x-a}{\sigma}\right)^{1/\gamma}\right)^{-1},&x>a\\ 0,&x\leq a\end{cases},$
where $10<\gamma<$ 1 and $\sigma>0$. Simple computations show that
$\mathbb{E}[|X-\mathbb{E}X|^{p}]=C(p,q,\gamma)\left(\mathbb{E}[|X-\mathbb{E}X|^{q}]\right)^{p/q}$
for $1<q<p<1/\gamma$, where
$I(p,\gamma)$ $:=\int_{0}^{+\infty}\left(x-\frac{\gamma\pi}{\sin\gamma\pi}\right)^{p}\frac{x^{1/\gamma}}{\gamma x(1+x^{1/\gamma})^{2}}\,{\rm d}x,$
$C(p,q,\gamma)$ $=I(p,\gamma)/I(q,\gamma)^{p/q}.$
However, the moment under this distribution is a function of $\sigma$; when $\sigma$ increases, the moment of any fixed order also increases, if it exists. The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment [5]. Therefore, for an arm with an unknown $\sigma$, we cannot apply the robust UCB policy. Furthermore, the learning efficiency can be improved if a larger $p$ can be used (tighter bound on the tail, but the corresponding moment bound becomes harder to know or compute a priori), it is desired to eliminate the need for any prior knowledge regarding the specific values or bounds of the moments. This motivates Lattimore [21] to consider a scale free algorithm. After we generalize Lattimore’s work, we will give a correction to a minor flaw in a lemma of Bubeck et al. [5] which subsequently caused a small error in the computation of regret upper bound in Lattimore [21].
1.2 Related Works
The theorems of the frequentist MAB have been developed for a long time and various milestones have been made since Lai and Robbins [20]. In contrast to the Bayesian model, the frequentist MAB does not assume a priori knowledge of the initial probability or state of the system but learns the core parameters solely through the past sampling history, e.g., computing an upper confidence bound as a function of the frequency that an arm is selected and the values of the observed samples. Such a way of decision-making in balancing exploration with exploitation contrasts to the Bayesian MAB where the tradeoff is usually addressed by dynamic programming for the time-varying characteristics of the system or the observation model [12]. After the UCB1 policy by Auer et al. in [3], many other policies of the UCB-type have been proposed in the literature. Maillard et al. [26] directed their attention to the prospect of KL divergence between probability distributions. They introduced the KL-UCB algorithm, which notably enhanced the regret bound. This improvement was achieved under the assumption of rewards being bounded and having a known upper bound. Furthermore, KL-UCB policy achieves the asymptotic lower bound of regret growth for this class of rewards given in Burnetas and Katehakis [6]. Kaufmann et al. [18] combined the Bayesian method with the UCB class and proposed Bayes-UCB for bounded rewards with a known bound, achieving a theoretical regret bound similar to that of KL-UCB. Numerical experiments demonstrated the high efficiency of learning by Bayes-UCB for Gaussian reward distributions. A survey of such UCB policies can be found in Burtini et al. [7]. These policies were neither implemented nor proved to achieve the logarithmic regret growth in the case of unbounded reward distributions. In 2011, the first author of this paper proposed UCB1-LT with a proven bound on the regret growth of the logarithmic order for the class of light-tailed reward distributions [24], filling the vacancy for the case of unbounded light-tailed reward distributions. Together with the extended robust UCB for the heavy-tailed class focused in this paper, we provide a relatively complete picture of UCB policies for general reward distributions. After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
For the case of heavy-tailed reward distributions, we have mentioned that Bubeck et al. [5] proposed the robust UCB policy under the assumption that an upper bound of the moments is known. The robust UCB policy uses different mean estimators rather than the empirical one and remarkably achieves the logarithmic regret growth rate. Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator [8]. The truncated mean estimator requires the knowledge of an upper bound of origin moments, which is not very effective in the case that the reward mean is far away from zero. The median-of-means estimator considers the (central) moment and thus patches up the deficiency of the truncated mean estimator: the mean estimation will not be affected when the reward distributions are changed by translations. The Catoni mean estimator is more complicated and can only be used when variances of the reward distributions exist ($p\geq 2$). But when applied in the robust UCB policy, it gives a much smaller regret coefficient than the other two estimators. Recently Chen et al. [11] showed a result of applying Catoni mean estimator to the case that $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators. Finer estimations when extending the Catoni mean estimator to the case of $p<2$ are interesting for future investigations. Several recent papers also considered the heavy-tailed MAB model and made significant progresses. Wei and Srivastava [36] proposed the robust MOSS policy with logarithmic regret growth. Agrawal et al. [1] established a policy called $\text{KL}_{\inf}$-UCB, which achieves a logarithmic upper bound of regret very close to the asymptotic lower bound proved by Burnetas and Katehakis [6]. For other non-classical MAB models, the heavy-tailed model has also been considered [37, 30]. However, all these policies do not discard the knowledge requirement of an upper bound of moments, which has been removed from the extended robust UCB policy proposed in this paper.
There are also policies not following the idea of UCB. The $\varepsilon$-greedy policy is well-studied, and many policies are proposed using its idea, e.g., the constant $\varepsilon$-decreasing policy [34], GreedyMix [9] and $\varepsilon_{n}$-greedy [3]. The Deterministic Sequencing of Exploration and Exploitation (DSEE) policy [25, 33] achieves the logarithmic regret for both cases of single player and multiple players (with the knowledge on moment upper bounds). Other recent progresses on variants of MAB, such as adversarial bandits, contextual bandits and linear bandits, can be found in Lattimore and Szepesvári [22, 13, 17].
2 The Heavy-Tailed Distribution
2.1 The Extended Robust UCB Policy
Consier $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$. In contrast with the light-tailed class, the heavy-tailed distribution may yield a very high (or low) reward realization with a high probability. Consequently, the empirical mean estimator $\overline{X}_{i,s}=\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for a sequence of i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work well as an estimation to the expectation of $X_{i}$ for arm $i$. The appendix of Bubeck et al. [5] has shown that under the assumption of heavy-tailed distributions, with probability at least $1-\delta$,
$\frac{1}{s}\sum_{j=1}^{s}X_{i,j}\leq\theta_{i}+\left(\frac{3\mathbb{E}|X_{i}-\theta_{i}|^{r}}{\delta s^{r-1}}\right)^{\frac{1}{r}},$
where $1<r\leq 2$. This bound is too loose to obtain a UCB policy achieving the logarithmic regret growth. To remedy this issue of divergence, we adopt the median-of-means estimator proposed in Bubeck et al. [5].
Definition 1.
For a finite i.i.d. sequence of random variables $\{X_{i,j}\}_{j=1}^{s}$ drawn from a distribution $F_{i}$, define the median-of-means estimator $\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)$ with $k$ $(k\leq s)$ bins as the median of $k$ empirical means
$\left\{\frac{1}{N}\sum_{j=lN+1}^{(l+1)N}X_{i,j}\right\}_{l=0}^{k-1},$
where $N=\lfloor s/k\rfloor$.
The following lemma offers an approximation of the proximity between the median-of-means estimator, utilizing a specific number of bins, and the actual mean. It should be noted that this lemma is similar to Lemma 2 in Bubeck et al. [5]. However, there is a minor flaw in the proof of the latter, which will be explained later. Moreover, there exist differences between the two lemmas, prompting us to restate and establish their distinctions in the following exposition.
Lemma 1.
Given an arbitrary $\varepsilon>0$. Suppose that
$\log\delta^{-1}>1/\varepsilon,\ s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\varepsilon\log(\delta^{-1})-1}$
and $v=\mathbb{E}[|X_{i}-\theta_{i}|^{p}]<\infty$, for some $p$ such that $p>1$. Then with probability at least $1-\delta$, the median-of-means estimator
$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\geq\theta_{i}-(12v)^{1/p}\left(\frac{(8+\varepsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
Also, with a probability of at least $1-\delta$,
$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\leq\theta_{i}+(12v)^{1/p}\left(\frac{(8+\varepsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
Proof.
Observe that $s\geq\lceil 8\log\delta^{-1}\rceil$ holds under the given condition in the lemma. For simplicity of presentation, we introduce some symbols as below. Let
$k$ $:=\lceil 8\log\delta^{-1}\rceil,$ (8)
$\chi$ $:=\frac{(8+\varepsilon)\log\delta^{-1}}{s},$ (9)
$\eta$ $:=(12v)^{1/p}\chi^{(p-1)/p},$ (10)
$\hat{\mu}_{l}$ $:=\frac{1}{N}\sum_{j=lN+1}^{(l+1)N}X_{i,j},\ \ l=0,1,\dots,k-1.$ (11)
According to the appendix of Bubeck et al. [5],
$\xi:=\mathbb{P}\left(\hat{\mu}_{l}>\theta_{i}+\eta\right)\leq\frac{3v}{N^{p-1}\eta^{p}}=\frac{1}{4N^{p-1}\chi^{p-1}}.$
Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$. Since $s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\varepsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$. Then using Hoeffding’s inequality for the tail of a binomial distribution [14], we have
$\mathbb{P}\left(\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)>\theta_{i}+\eta\right)\leq\mathbb{P}\left(\sum_{l=0}^{k-1}\mathbb{I}(\hat{\mu}_{l}\geq\theta_{i}+\eta)\geq\frac{k}{2}\right)\leq\exp(-2k(1/2-\xi)^{2})\leq\delta.$
Similarly, the other inequality also holds by symmetry. ∎
With the help of the median-of-means estimator, we can now define the upper confidence bound for the extended robust UCB policy. As mentioned above, we assume that for any arm distribution $X_{i}\in\mathcal{F}$ with mean $\theta_{i}$, there is a moment control coefficient $C_{p,q}$ ($1<q<p$) such that
$\mathbb{E}[|X_{i}-\theta_{i}|^{p}]\leq C_{p,q}\mathbb{E}[|X_{i}-\theta_{i}|^{q}]^{r},$ (12)
where $r=p/q$.
Definition 2.
Recall the notations introduced by (8) and (9). Given an arbitrary $\varepsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
$\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta):=$
[…]
where
$C^{\prime}=\left(12(C_{p,q}+1)\right)^{q/p}.$ (13)
Note that this upper confidence bound always exists, since the left-hand side of the above inequality tends to $\infty$- ∞ as $\theta\to-\infty$, while the right-hand side tends to +$\infty$+ ∞.
Lemma 2.
Suppose that the conditions in Lemma 1 are satisfied, then with probability at least $1-2\delta$, we have
$\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)\geq\theta_{i}.$
Proof.
Recall the notations introduced by (8), (9) and (13). If $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)=+\infty$, the inequality holds with probability $1$. Assume that $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)<+\infty$, i.e.
$01-C^{\prime}\chi^{(p-q)/p}>$ 0 .
Define
$Y_{i}$ $:=|X_{i}-\theta_{i}|^{q},$
$Y_{i,j}$ $:=|X_{i,j}-\theta_{i}|^{q},$ (14)
$v_{q}$ $:=\mathbb{E}Y_{i},$
$v_{p}$ $:=\mathbb{E}[|X_{i}-\theta_{i}|^{p}].$
From Lemma 1, we have
[…]
From Lemma 1 and by the definition of $C_{p,q}$, we have
$\begin{split}\mathbb{E}[\left|Y_{i}-v_{q}\right|^{p/q}]&\leq\mathbb{E}[Y_{i}^{p/q}+v_{q}^{p/q}]=\mathbb{E}\left[|X_{i}-\theta_{i}|^{p}+v_{q}^{p/q}\right]\\ &=v_{p}+v_{q}^{p/q}\leq(C_{p,q}+1)v_{q}^{p/q}.\end{split}$ (16)
Merging the two inequalities (15) and (16), with probability at least $1-2\delta$, we have
$\theta_{i}\leq\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)+\left(\frac{12\hat{\mu}(\{|X_{i,j}-\theta_{i}|^{q}\}_{j=1}^{s},k)}{\max\left\{0,1-C^{\prime}\chi^{(p-q)/p}\right\}}\right)^{1/q}\chi^{(q-1)/q}.$
The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. ∎
Corollary 1.
Under the conditions of Lemma 1, with probability at least $1-\delta$, we have
$\hat{\mu}(\{Y_{i,j}\}_{j=1}^{s},k)-v_{q}\leq C^{\prime}v_{q}\chi^{(p-q)/p}.$
Proof.
This corollary is a direct consequence of Lemma 1 and the inequality in (16). ∎
Finally, with the upper confidence bound estimator $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s})$, we complete the extended robust UCB policy as follows.
1 Input: $\varepsilon>0$, $K$: the number of arms
2 Initialize: $s_{i}\leftarrow 0$ and $\{X_{i,j}\}_{j=1}^{s_{i}}$ an empty sequence for each arm $i$, $1\leq i\leq K$
3 for $t\leftarrow 1$ to $T$ do
4       for $i\leftarrow 1$ to $K$ do
5             if $s_{i}=0$ then
6                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
7             end if
8            else
9                  Compute arm $i$’s upper confidence bound as $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$
10             end if
11
12       end for
13      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
14       Update: $s_{j}\leftarrow s_{j}+1$ and append $x$ to sequence $\left\{X_{j,k}\right\}_{k=1}^{s_{j}}$
15 end for
Algorithm 1 The Extended Robust UCB Policy
Here we present the main theorem that the extended robust UCB policy achieves the logarithmic order of regret growth.
Theorem 1.
Suppose $p>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$, $1<q<p$ and $C_{p,q}$ such that (12) holds. Then for any $\varepsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
To prove Theorem 1, we establish the following core lemma to bound the time $s_{i}$ spent on each arm $i$ by the extended robust UCB policy.
Lemma 3.
The expected number of times $\mathbb{E}[s_{i}(T)]$ that a sub-optimal arm is chosen by the extended robust UCB has an upper bound of the logarithmic order with $T$.
Proof.
In addition to notations (8), (9) and (13), we introduce two new notations:
$\tau(x)$ $:=\frac{1}{1-C^{\prime}x^{(p-q)/p}},$ (17)
$\zeta$ $:=12\tau(\chi)\hat{\mu}\left(\{|X_{i,j}-\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)|^{q}\},k\right).$ (18)
By the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$, we have
$\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)=\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)+\zeta^{1/q}\chi^{(q-1)/q}.$ (19)
We need to consider the probability of the following event:
$\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)-\theta_{i}\leq\Delta_{i},$ (20)
where $\Delta_{i}$ is defined in (2). First we estimate an upper bound of $\zeta$. Suppose that $\tau(\chi)>0$, we have
$\left|X_{i,j}-\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)\right|^{q}\leq 2^{q-1}|X_{i,j}-\theta_{i}|^{q}+2^{q-1}|\theta_{i}-\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)|^{q}.$
Recall that $Y_{i,j}=|X_{i,j}-\theta_{i}|^{q}$ defined in (14). Thus
$\zeta\leq$ $12\tau(\chi)\left(2^{q-1}\hat{\mu}\left(\{Y_{i,j}\}_{j=1}^{s},k\right)+2^{q-1}|\theta_{i}-\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)|^{q}\right)$
$\leq$ $12\tau(\chi)\left(2^{q-1}\hat{\mu}\left(\{Y_{i,j}\}_{j=1}^{s},k\right)\right.$
$\left.+2^{2q-2}\left(|\theta_{i}-\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)|^{q}+|\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)-\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)|^{q}\right)\right)$
$=$ $12\tau(\chi)\left(2^{q-1}\hat{\mu}\left(\{Y_{i,j}\}_{j=1}^{s},k\right)+2^{2q-2}\left(|\theta_{i}-\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)|^{q}+\zeta\chi^{q-1}\right)\right).$
The inequality above uses (19) and also the fact that
$|a+b|^{q}\leq 2^{q-1}(|a|^{q}+|b|^{q})$
by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality [29]. Suppose that
$1>3\cdot 2^{2q}\chi^{q-1}\tau(\chi),$ (21)
then
$\zeta\leq\frac{12\tau(\chi)\left(2^{q-1}\hat{\mu}(\{Y_{i,j}\}_{j=1}^{s},k)+2^{2q-2}|\theta_{i}-\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)|^{q}\right)}{1-3\cdot 2^{2q}\chi^{q-1}\tau(\chi)}.$
Suppose that $\delta$ and $s$ meet the conditions in Lemma 1. Then with probability at least $1-2\delta$, we have
$|\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)-\theta_{i}|\leq(12v_{q})^{1/q}\chi^{(q-1)/q}.$ (22)
Using (22) and Corollary 1, by defining a function $B(x)$ for simplicity as
$B(x):=\frac{6\cdot 2^{q}\tau(x)}{1-3\cdot 2^{2q}x^{q-1}\tau(x)}\left(1+C^{\prime}x^{(p-q)/p}\right)+\frac{3\cdot 2^{2q}\tau(x)\cdot 12x^{(q-1)}}{1-3\cdot 2^{2q}x^{q-1}\tau(x)},$ (23)
then with probability at least $1-3\delta$, we have
$\zeta\leq v_{q}B(\chi).$ (24)
Now we go back to the upper bound estimation of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)-\theta_{i}$. Using Lemma 1 again and (24), with probability at least $1-3\delta$, we have
$\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)-\theta_{i}=$ $(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)-\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k))+(\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)-\theta_{i})$
$=$ $\zeta^{1/q}\chi^{(q-1)/q}+(\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)-\theta_{i})$
$\leq$ $v_{q}^{1/q}B(\chi)^{1/q}\chi^{(q-1)/q}+(12v_{q})^{1/q}\chi^{(q-1)/q}$
$=$ $v_{q}^{1/q}\chi^{(q-1)/q}\left(B(\chi)^{1/q}+12^{1/q}\right).$
Therefore, if
$v_{q}^{1/q}\chi^{(q-1)/q}\left(B(\chi)^{1/q}+12^{1/q}\right)\leq\Delta_{i}$ (25)
holds, then (20) is true with probability at least $1-3\delta$.
Again, we need to emphasize the assumptions under which (20) is true with probability at least $1-3\delta$: $\tau(\chi)>0$, (21) and (25). Equivalently, the following three inequalities are required for $x=\chi$:
$1>$ $C^{\prime}x^{(p-q)/p},$ (26)
$1>$ $3\cdot 2^{2q}x^{q-1}\tau(x),$ (27)
$\Delta_{i}\geq$ $v_{q}^{1/q}x^{(q-1)/q}(B(x)^{1/q}+12^{1/q}).$ (28)
[…]
$\lim_{x\searrow 0}x^{(q-1)/q}(B(x)^{1/q}+12^{1/q})=0.$
So one can always choose a $\chi_{M}>0$ such that (26), (27) and (28) hold for $x\leq\chi_{M}$.
Note that we also need the conditions $\log\delta^{-1}>1/\varepsilon$ and
$s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log\delta^{-1}+1}{\varepsilon\log\delta^{-1}-1}$
to apply Lemma 1.
Next, we bound $\mathbb{E}[s_{i}(T)]$ for any $i$ such that $\theta_{i}<\theta_{1}$ and prove the logarithmic regret growth by (1). Define
$\mathbb{I}\left(X\right):=\begin{cases}1,&X\text{ occurs}\\ 0,&X\text{ doesn't occur}\end{cases}$
to denote the characteristic function of an event $X$. Let $\delta=t^{-2}$. We use the notation $A_{t}=i$ to indicate that arm $i$ is chosen at time $t$. For any $i$ such that $\theta_{i}\neq\theta_{1}$, note that $A_{t}=i$ implies $\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})$. We have
$\begin{split}&s_{i}(T)=\sum_{t=1}^{T}\mathbb{I}\left(A_{t}=i\right)\\ \leq&\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)+\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})\geq\theta_{1}\right).\end{split}$ (29)
We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\varepsilon$ is equivalent to $t>\exp\left(\frac{1}{2\varepsilon}\right)$. Write
$l_{0}:=\max\left\{\left\lfloor\exp\left(\frac{1}{2\varepsilon}\right)\right\rfloor,\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil\right\}.$ (30)
Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\varepsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
$\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)$
$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right)$
$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\varepsilon)\log t^{2}\frac{8\log t^{2}+1}{\varepsilon\log t^{2}-1}\right).$
By Lemma 2, we have
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)\right]\leq l_{0}+\sum_{t=1}^{\infty}\frac{2}{t^{2}}\leq l_{0}+\frac{\pi^{2}}{3}.$
Now we bound the second part of the right hand side of (LABEL:s_i_divide). Define
$l_{1}:=\max\left\{l_{0},\left\lceil\frac{8+\varepsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}.$ (31)
For sufficiently large $T$, we have
$\begin{split}&\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})\geq\theta_{1}\right)\\ \leq&l_{1}+\sum_{t=l_{1}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})-\theta_{i}\geq\Delta_{i}\text{ and }s_{i}(t-1)\geq l_{1}\right).\end{split}$ (32)
By the definition of $\chi$ in (9) and the choice of $\chi_{M}$, using (20), we have
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})\geq\theta_{1}\right)\right]\leq l_{1}+\frac{\pi^{2}}{2}.$
Finally, we have
$\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}.$
For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\varepsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (33)
A Monte Carlo simulation example based the Pareto distribution of type III (see Sec. 1.1) is shown in Fig. 1. Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e., $\chi$ defined in (9) becomes sufficiently small such that, in the $\sup$ expression of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is nonzero and the growth rate of the right-hand side with $\theta$ becomes small).
Figure 1: The performance of extended robust UCB
2.2 Further Analysis and Comparison of the Extended Robust UCB
Now we consider scenarios where specific values of $p$ and $q$ are chosen to further improve the efficiency of the extended robust UCB policy.
Case 1. Suppose that the $p^{\prime}$-th order of moments of reward distributions exist with $p^{\prime}>2$ for all arms. Then we can choose $p$, $q$ such that $p=2q\leq p^{\prime}$. The inequality (16) can be refined to
$\begin{split}\mathbb{E}[|Y_{i}-v_{q}|^{p/q}]&=\mathbb{E}Y_{i}^{2}+v_{q}^{2}-2\mathbb{E}v_{q}Y_{i}=\mathbb{E}Y_{i}^{2}-v_{q}^{2}=v_{p}-v_{q}^{2}\\ &\leq(C_{p,q}-1)v_{q}^{p/q}.\end{split}$ (34)
Hence, we can refine $C^{\prime}$ from (13) to
$C^{\prime}=\sqrt{12(C_{p,q}-1)},$ (35)
which can be much smaller than the original definition (13).
Let’s come back to the definition of $\chi_{M}$, which guarantees the three inequalities (26), (27) and (28) to hold. The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e., the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy may converge to the best arm faster.
Case 2. In addition to the assumption in Case 1, we add an assumption that the $4$-th order moments of reward distributions exist as in Lattimore [21]. Thus we can choose $p=4$ and $q=2$. In this case, our UCB in Definition 2 can be rewritten as
[…]
In this case $C_{p,q}$ degenerates to the upper bound on kurtosis. The function $B(x)$ defined in (23) becomes
$B(x)=\frac{24\tau(x)}{1-48x\tau(x)}\left(1+C^{\prime}\sqrt{x}\right)+\frac{576\sqrt{x}\tau(x)}{1-48x\tau(x)},$
and the three inequalities (26), (27) and (28) are refined to
11 11 $>C^{\prime}\sqrt{x},$ (36)
11 11 $>48x\tau(x),$ (37)
$\Delta_{i}$ $\geq\sqrt{v_{2}x}\left(\sqrt{B(x)}+2\sqrt{3}\right).$ (38)
Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (33) will be significantly decreased. Now we compare the extended robust UCB and the scale free algorithm proposed in Lattimore [21], under the same assumptions specified at the beginning of Case 2.
We first address a flaw within Lemma 2 of Bubeck et al. [5]. This lemma, crucial in Lattimore’s analysis [21] for deriving the regret upper bound of the scale free algorithm, is examined more closely below. For clarity and coherence with the notations employed in this paper, we present the following adapted version.
A Flawed Statement [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $\mathbb{E}|X_{i}-\theta_{i}|^{p}=v_{p}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
$\hat{\mu}\left(\{X_{i,j}\}_{j=1}^{s},k^{\prime}\right)\leq\theta_{i}+(12v_{p})^{1/p}\left(\frac{16\log({\rm e}^{1/8}\delta^{-1})}{s}\right)^{(p-1)/p}.$
The last inequality in the proof of this statement in Bubeck et al. [5] assumes that $\exp(-k^{\prime}/8)\leq\delta$, and this is equivalent to $k^{\prime}\geq 8\log\delta^{-1}$. However, the definition of $k^{\prime}$ in this statement is
$k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor.$
To ensure that the inequality $\exp(-k^{\prime}/8)\leq\delta$ in their proof to hold, it is required that
$\lfloor s/2\rfloor\geq 8\log\delta^{-1},$ (39)
which is not assumed in the statement. As the statement is applied to the robust UCB, under the choice $\delta=t^{-2}$ and the definition $s:=s_{i}(t-1)$ as the number of times that arm $i$ was chosen before time $t$, Assumption (39) becomes
$\lfloor s_{i}(t-1)/2\rfloor\geq 16\log t.$
Hence the regret upper bound in Theorem 3 of Bubeck et al. [5] is at least $32\Delta_{i}\log T$.
Below we give the corrected version of the above statement. Again, we will adopt notations introduced in this paper for consistency.
Lemma 4.
Let $\delta\in(0,1)$ and $p\in(1,2]$. Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d. random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $\mathbb{E}|X_{i}-\theta_{i}|^{p}=v_{p}$, where $s$ is chosen such that $\lfloor s/2\rfloor\geq 8\log\delta^{-1}$. Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$. Then with probability at least $1-\delta$,
$\hat{\mu}\left(\{X_{i,j}\}_{j=1}^{s},k^{\prime}\right)\leq\theta_{i}+(12v_{p})^{1/p}\left(\frac{16\log({\rm e}^{1/8}\delta^{-1})}{s}\right)^{(p-1)/p}.$
Now we look into the results in Lattimore [21], whose Lemma 4 is a special version of the flawed Lemma 2 in Bubeck et al. [5]. Lattimore’s Lemmas 5 and 6 in [21] both use the flawed version to estimate the regret upper bound as in the proof of their Theorem 2. We do not intend to give details about fixing all of them, but it is important to point out that, after applying the correct version (i.e., adding Assumption (39)) as in Lemma 4, the upper bound estimation of $\mathbb{E}[s_{i}(T)]$ in Lattimore [21] is at least (the constants introduced in their paper are explicitly calculated with notations translated here):
$\begin{split}3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}\log({\rm e}^{1/8}\delta^{-1})+16\log\delta^{-1}.\end{split}$ (40)
The next theorem shows that for properly chosen parameter $\varepsilon>0$, the upper bound given in (33) is also tighter than (40) over long-run if $\Delta_{i}$ is small, i.e., arms are hard to be distinguished (the low-discrimination case).
Theorem 2.
Choose any $8<\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$ ($\approx 44.158$) and let $\delta=t^{-2}$. The upper bound given in (40) of $\mathbb{E}[s_{i}(T)]$ is greater than that given in (33) for sufficiently large $T$ and sufficiently small $\Delta_{i}$. That is, by dividing both the upper bound estimations by $\log\delta^{-1}$ and letting $T\to+\infty$, we have
$(8+\varepsilon)\left(\frac{8}{\varepsilon}+\max\left\{\frac{8}{\varepsilon},\frac{1}{\chi_{M}}\right\}\right)<3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}+16$ (41)
for sufficiently small $\Delta_{i}$. The regret upper bound of the scale free algorithm proposed in Lattimore [21] is thus larger than that of the extended robust UCB policy for $\varepsilon\in(8,\frac{280-16\sqrt{2}}{2\sqrt{2}+3})$.
Proof.
First note that, since $\Delta_{i}$ is sufficiently small and $\varepsilon$, $C^{\prime}$ are not related to $\Delta_{i}$, we only need to show (noticing that $\varepsilon>8$ implies $(8+\varepsilon)\frac{8}{\varepsilon}\leq 16$)
$(8+\varepsilon)\frac{1}{\chi_{M}}<3648\frac{v_{2}}{\Delta_{i}^{2}}.$ (42)
Equivalently, it is sufficient to show that
$x^{\prime}:=\frac{(8+\varepsilon)\Delta_{i}^{2}}{3648v_{2}}<\chi_{M}$ (43)
for sufficiently small $\Delta_{i}$. In order to prove the above, we need to show that (36), (37) and (38) are all true for $x=x^{\prime}$ based on the definition of $\chi_{M}$. For (36), let $x=x^{\prime}$, we have $1>\frac{C^{\prime}\Delta_{i}}{8}\sqrt{\frac{\varepsilon+8}{57v_{2}}}$, which is true for small $\Delta_{i}$. Furthermore, (37) becomes
$1>6(8+\varepsilon)\Delta_{i}^{2}\frac{1}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(8+\varepsilon)v_{2}}},$
which also holds for sufficiently small $\Delta_{i}$. For (38), let $x=x^{\prime}$ and the inequality becomes
$(\varepsilon+8)\left(\sqrt{6}\sqrt{\frac{v_{2}\left(\sqrt{57}(C^{\prime}+24)\Delta_{i}\sqrt{\frac{\varepsilon+8}{v_{2}}}+456\right)}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(\varepsilon+8)v_{2}}-6\Delta_{i}^{2}(\varepsilon+8)}}+\sqrt{3}\right)^{2}<912.$ (44)
Taking the limit $\Delta_{i}\to 0$, (44) becomes $\left(6\sqrt{2}+9\right)(\varepsilon+8)<912$, which holds for $\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. By continuity, we conclude that (43) holds for sufficiently small $\Delta_{i}$ if $\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. ∎
Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound by only a constant factor and a constant term as in the scale free algorithm of Lattimore [21] (see (40)). First, we restate the lower bound derived in Lattimore [21] as follows.
Theorem 3.
Let $H_{\kappa_{0}}$ be the set of distributions that has kurtosis less than $\kappa_{0}$. Assuming $\Delta>0$ and $\kappa_{0}\geq 7/2$. Suppose that $X\in H_{\kappa_{0}}$ has a mean of $\mu$, a positive variance of $\sigma^{2}$, and a kurtosis of $k$. Then
$\begin{split}\inf\{{\rm KL}(X,X^{\prime}):&X^{\prime}\in H_{\kappa_{0}}\ \text{and}\ \mathbb{E}[X^{\prime}]>\mu+\Delta\}\\ \leq&\begin{cases}\min\{-\log(1-p),\frac{C_{1}\Delta^{2}}{\sigma^{2}}\},&\text{if}\ C_{0}\sqrt{k}(k+1)\frac{\Delta}{\sigma}<\kappa_{0}\\ -\log(1-p),&\text{otherwise}\end{cases},\end{split}$ (45)
[…]
Based on this, we can draw the following conclusion:
Theorem 4.
Let $\delta=t^{-2}$. Under case 2, the regret of the extended robust UCB policy applied on $H_{\kappa_{0}}$ ($\kappa_{0}\geq 7/2$) differs from the lower bound of the regret by a constant factor and a constant term.
Proof.
If $\Delta_{i}$ of arm $i$ is large, that is, if (36) and (37) directly lead to (38), then $\mathbb{E}[s_{i}(T)]/\log T$ is upper bounded by a constant not related to $\Delta_{i}$. We are interested in the case where $\Delta_{i}$ is small, i.e., when (38) becomes the main restraint of $\chi_{M}$:
$\Delta_{i}=\sqrt{v_{2}\chi_{M}}\left(\sqrt{B(\chi_{M})}+2\sqrt{3}\right).$ (46)
By (36), since $B(x)$ is an increasing function of $x$, $B(\chi_{M})$ is bounded by a constant only related to $C^{\prime}$. We have
$\frac{1}{\chi_{M}}\leq C_{B}\frac{v_{2}}{\Delta_{i}^{2}},$ (47)
where $C_{B}=B(1/{C^{\prime}}^{2})$ is a constant. Note that
$\begin{split}\lim\sup\frac{\mathbb{E}[s_{i}(T)]}{\log T}&\leq 2(8+\varepsilon)\left(1+\max\left\{\frac{8}{\varepsilon},\frac{1}{\chi_{M}}\right\}\right)\\ &\leq 2(8+\varepsilon)\left(1+\max\left\{\frac{8}{\varepsilon},C_{B}\frac{v_{2}}{\Delta_{i}^{2}}\right\}\right).\end{split}$ (48)
With the conclusion of (4) and Theorem 3, we finished the proof. ∎
3 Light-Tailed Reward Disdtributions
In this section, we formally and briefly discuss the past work of the first author of this paper for light-tailed reward distributions [24] and its comparisons with some subsequent work following this line.
3.1 The UCB1-LT Policy
The empirical mean is adopted in estimating the upper confidence bound for the class of light-tailed reward distributions. As before, let $\overline{X}_{i,s}$ denote the empirical mean $\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for arm $i$, where $\{X_{i,j}\}_{j=1}^{s}$ forms the i.i.d. random reward sequence drawn from an unknown distribution $X_{i}\in\mathcal{F}$.
Lemma 5.
(Bernstein-type bound) For i.i.d. random variables $\{X_{i,j}\}_{j=1}^{s}$ drawn from a light-tailed distribution $X_{i}$ (with mean $\theta_{i}$) with a finite moment-generating function $M(u)$ over range $u\in[-u_{0},u_{0}]$ for some $u_{0}>0$. We have, $\forall\varepsilon>0$,
$\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\geq\varepsilon)\leq\begin{cases}\exp(-\frac{s}{2\zeta}\varepsilon^{2}),&\varepsilon<\zeta u_{0}\\ \exp(-\frac{su_{0}}{2}\varepsilon),&\varepsilon\geq\zeta u_{0}\\ \end{cases},$ (49)
where $\zeta>0$ satisfies $\zeta\geq\sup_{|u|\leq u_{0}}{M^{(2)}(u)}$. A similar bound holds for $\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\leq-\varepsilon)$ by symmetry.
The proof of the above lemma follows a similar argument as in Vershynin [35]. Using the Bernstein-type bound by Lemma 5, we propose the UCB1-LT policy as follows.
1Input: $a_{1}\geq 8\zeta$, $a_{2}\geq a_{1}/(\zeta u_{0})$, $K$: the number of arms
2 Initialize: $s_{i}\leftarrow 0$ and $\overline{X}_{i}\leftarrow 0$ for each arm $i$, $1\leq i\leq K$
3 for $t\leftarrow 1$ to $T$ do
4       for $i\leftarrow 1$ to $K$ do
5             if $s_{i}=0$ then
6                  Assign +$\infty$+ ∞ to the upper confidence bound
7             end if
8            else
9                   if $\sqrt{\frac{a_{1}\log t}{s_{i}}}<\zeta u_{0}$ then
10                         Compute the upper confidence bound as $\overline{X}_{i}+\sqrt{\frac{a_{1}\log t}{s_{i}}}$
11                   end if
12                  else
13                         Compute the upper confidence bound as $\overline{X}_{i}+\frac{a_{2}\log t}{s_{i}}$
14                   end if
15
16             end if
17
18       end for
19      Choose arm $j$ that maximizes the upper confidence bound and obtain reward $x$
20       $\overline{X}_{j}\leftarrow\frac{x+s_{j}\overline{X}_{j}}{s_{j}+1}$ and $s_{j}\leftarrow s_{j}+1$
21 end for
Algorithm 2 The UCB1-LT Policy
The UCB1-LT policy considers two upper confidence bounds and alternatively uses one of them according to values of $t$ and $s_{i}(t-1)$. To minimize the theoretical regret upper bound, we can choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$. The following theorem shows that UCB1-LT achieves the logarithmic regret growth for light-tailed reward distributions.
Theorem 5.
For light-tailed reward distributions, the regret of the UCB1-LT policy satisfies the following inequality:
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{i:\theta_{i}<\theta_{1}}\Delta_{i}\left(\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}\log T+1+\frac{\pi^{2}}{3}\right),$ (50)
Proof.
Define
$c(t,s):=\begin{cases}\sqrt{\frac{a_{1}\log t}{s}},&\sqrt{\frac{a_{1}\log t}{s}}<\zeta u_{0}\\ \frac{a_{2}\log t}{s},&\sqrt{\frac{a_{1}\log t}{s}}\geq\zeta u_{0}\end{cases}.$
Similar to the procedure in Auer et al. [3], for any integer $L>0$ and $n$ such that $\theta_{i}<\theta_{1}$, we have
$\mathbb{E}[s_{i}(T)]\leq$ $L+\sum_{t=1}^{T}\mathbb{P}\{\overline{X}_{i,t}+c(t,s_{i}(t-1))\geq\overline{X}_{1,t}+c(t,s_{1}(t-1))\ \text{and}\ s_{i}(t-1)>L\}$
$\leq$ $L+\sum_{t=1}^{\infty}\sum_{s=1}^{t-1}\sum_{k=L}^{t-1}\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}+c(t,k)\geq\frac{1}{s}\sum_{j=1}^{s}X_{1,j}+c(t,s)\right)$
$\leq$ $L+\sum_{t=1}^{\infty}\sum_{s=1}^{t-1}\sum_{k=L}^{t-1}\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}\geq\theta_{i}+c(t,k)\right)$
$+\mathbb{P}\left(\frac{1}{s}\sum_{j=1}^{s}X_{1,j}\leq\theta_{1}-c(t,s)\right)+\mathbb{P}\left(\theta_{i}+2c(t,k)>\theta_{1}\right).$
Choose $L_{0}=\lceil\max\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\}\rceil$, $\forall k\geq L_{0}$, we have
$c(t,k)\leq$ $\max\left\{\sqrt{\frac{a_{1}\log t}{k}},\frac{a_{2}\log t}{k}\right\}\leq\max\left\{\sqrt{\frac{a_{1}\log t}{L_{0}}},\frac{a_{2}\log t}{L_{0}}\right\}$
$\leq$ $\max\left\{\sqrt{a_{1}\log t\frac{(\theta_{1}-\theta_{i})^{2}}{4a_{1}\log T}},a_{2}\log t\frac{\theta_{1}-\theta_{i}}{2a_{2}\log T}\right\}$
$\leq$ $\frac{\theta_{1}-\theta_{i}}{2}.$
Then
$\mathbb{E}[s_{i}(T)]\leq L_{0}+\sum_{t=1}^{\infty}\sum_{s=1}^{t-1}\sum_{k=L_{0}}^{t-1}\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}\geq\theta_{i}+c(t,k)\right)$
$+\mathbb{P}\left(\frac{1}{s}\sum_{j=1}^{s}X_{1,j}\leq\theta_{1}-c(t,s)\right).$
Now we bound the probabilities by the Bernstein-type bound (49). If
$\sqrt{\frac{a_{1}\log t}{k}}<\zeta u_{0},$
then
$\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}\geq\theta_{i}+c(t,k)\right)=$ $\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}\geq\theta_{i}+\sqrt{\frac{a_{1}\log t}{k}}\right)$
$\leq\exp\left(-\frac{k}{2\zeta}\left(\sqrt{\frac{a_{1}\log t}{k}}\right)^{2}\right)\leq$ $t^{-4};$
otherwise
$\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}\geq\theta_{i}+c(t,k)\right)=$ $\mathbb{P}\left(\frac{1}{k}\sum_{j=1}^{k}X_{i,j}\geq\theta_{i}+\frac{a_{2}\log t}{k}\right)$
$\leq\exp\left(-\frac{ku_{0}}{2}\frac{a_{2}\log t}{k}\right)\leq$ $t^{-4}.$
The same argument also applies to $\mathbb{P}\left(\frac{1}{s}\sum_{j=1}^{s}X_{1,j}\leq\theta_{1}-c(t,s)\right)$. Together we have
$\mathbb{E}[s_{i}(T)]\leq L_{0}+2\sum_{t=1}^{\infty}\sum_{s=1}^{t-1}\sum_{k=L_{0}}^{t-1}t^{-4}\leq\max\left\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\right\}+1+\frac{\pi^{2}}{3},$
A direct substitution of the above into (1) completes the proof. ∎
3.2 From UCB1-LT to $(\alpha,\psi)$-UCB
After the establishment of UCB1-LT by Liu and Zhao [24], Bubeck and Cesa-Bianchi [4] subsequently proposed $(\alpha,\psi)$-UCB under a more general assumption on the moment generating functions. Specifically, the $(\alpha,\psi)$-UCB policy assumes that the distribution of reward $X_{i}\in\mathcal{F}$ satisfies the following condition: there exists a convex function $\psi$ defined on $\mathbb{R}$ such that for all $\lambda>0$,
$\log\mathbb{E}\left[{\rm e}^{\lambda(X_{i}-\theta_{i})}\right]\leq\psi(\lambda)\text{ and }\log\mathbb{E}\left[{\rm e}^{\lambda(\theta_{i}-X_{i})}\right]\leq\psi(\lambda).$ (51)
For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (51) becomes the well-known Hoeffding’s lemma [4]. Since +$\infty$+ ∞ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT. By using the Legendre-Frenchel transform $\psi^{*}$ of $\psi$, defined as
$\psi^{*}(\varepsilon):=\sup_{\lambda\in\mathbb{R}}\lambda\varepsilon-\psi(\lambda).$
The upper confidence bound of $(\alpha,\psi)$-UCB at time $t$ is defined as
$I_{i}(t):=\overline{X}_{i,s_{i}(t-1)}+(\psi^{*})^{-1}\left(\frac{\alpha\log t}{s_{i}(t-1)}\right).$ (52)
Theorem 6 ([4]).
Assume that the reward distributions satisfy (51). Then $(\alpha,\psi)$-UCB with $\alpha>2$ achieves
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{\theta_{i}<\theta_{1}}\left(\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\log T+\frac{\alpha}{\alpha-2}\right).$ (53)
For the UCB1-LT policy, we can choose
$\psi(\lambda)=\begin{cases}\frac{\zeta\lambda^{2}}{2},&\lambda\leq u_{0}\\ +\infty,&\lambda>u_{0}\end{cases}.$ (54)
Then (51) becomes the same assumption as in UCB1-LT (see (6)). In this case, we have
$\psi^{*}(\varepsilon)=\begin{cases}\frac{\varepsilon^{2}}{2\zeta},&|\varepsilon|<\zeta u_{0}\\ u_{0}(|\varepsilon|-\frac{\zeta u_{0}}{2}),&|\varepsilon|\geq\zeta u_{0}\end{cases}.$
Thus the inverse $(\psi^{*})^{-1}(x)$ of $\psi^{*}$ for $x\geq 0$ will be
$(\psi^{*})^{-1}(x)=\begin{cases}\sqrt{2\zeta x},&0\leq x<\frac{\zeta u_{0}^{2}}{2}\\ \frac{x}{u_{0}}+\frac{\zeta u_{0}}{2},&x\geq\frac{\zeta u_{0}^{2}}{2}\end{cases}.$
Furthermore, if we choose $\alpha=4$, then the upper confidence bound defined in (52) becomes
$I_{i}(t)=\overline{X}_{i,s_{i}(t-1)}+\begin{cases}\sqrt{\frac{8\zeta\log t}{s_{i}(t-1)}},&\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}\\ \frac{4\log t}{u_{0}s_{i}(t-1)}+\frac{\zeta u_{0}}{2},&\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}\end{cases}.$ (55)
Notice that, if we choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$ in UCB1-LT, we get the same index function under the condition $\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}$. For the case $\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}$, the upper confidence bound of $(\alpha,\psi)$-UCB is not larger than that of UCB1-LT and thus may achieve a smaller theoretical upper bound of regret as follows:
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{\theta_{i}<\theta_{1}}\left(r(\Delta_{i})\log T+2\right),$ (56)
where
$r(\Delta_{i})=\begin{cases}\frac{32\zeta}{\Delta_{i}},&\Delta_{i}<2\zeta u_{0}\\ \frac{8\Delta_{i}}{u_{0}(\Delta_{i}-\zeta u_{0})},&\Delta_{i}\geq 2\zeta u_{0}\end{cases}.$
Note that this regret bound is better than (50) in Theorem 5 for UCB1-LT as $T$ becomes sufficiently large, as shown in the following lemma.
Lemma 6.
Choose the parameter $\alpha$ in $(\alpha,\psi)$-UCB such that $\alpha\leq 4$, and assume that Assumption (6) is satisfied and thus $\psi(\lambda)$ is chosen as in (54). Then the logarithmic regret upper bound in Theorem 6 has a leading constant no larger than that in Theorem 5, i.e.,
$\sum_{\theta_{i}<\theta_{1}}\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\leq\sum_{\theta_{i}<\theta_{1}}\Delta_{i}\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}.$
Proof.
We only need to show that
$\frac{\alpha x}{\psi^{*}(x/2)}\leq x\max\left\{\frac{4a_{1}}{x^{2}},\frac{2a_{2}}{x}\right\}$ (57)
for any $x>0$. From (54), we have
$\frac{\alpha x}{\psi^{*}(x/2)}=\begin{cases}\frac{8\alpha\zeta}{x},&0<x<2\zeta u_{0}\\ \frac{2\alpha x}{u_{0}(x-\zeta u_{0})},&x\geq 2\zeta u_{0}\end{cases}.$
From UCB1-LT, we have $a_{1}\geq 8\zeta$ and $a_{2}\geq a_{1}/(\zeta u_{0})\geq 8/u_{0}$, then
$x\max\left\{\frac{4a_{1}}{x^{2}},\frac{2a_{2}}{x}\right\}\geq\max\left\{\frac{32\zeta}{x},\frac{16}{u_{0}}\right\}=\begin{cases}\frac{32\zeta}{x},&0<x<2\zeta u_{0}\\ \frac{16}{u_{0}},&x\geq 2\zeta u_{0}\end{cases}.$
Since $\frac{8\alpha\zeta}{x}\leq\frac{32\zeta}{x}$ as $\alpha\leq 4$ and $\frac{2\alpha x}{u_{0}(x-\zeta u_{0})}\leq\frac{16}{u_{0}}$ as $x\geq 2\zeta u_{0}$, (57) is proved. ∎
Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments. Assume that the arms are offering rewards according to $19$ normal distributions. The means of reward distributions are $0.5,1.0,1.5,\dots,9.5$ with standard deviations all set to $60$. A Monte Carlo (MC) simulation with 100 runs and time period $T=50000$ is shown below. The horizontal axis in both figures is the time step $t$. The first figure only shows $T$ from $10000$ to $50000$ for better display, while the second one ranges from $1$ up to $50000$. The vertical axis of the left figure is the logarithmic regret averaged from the 100 MC runs (to approximate the expectation), while of the right one is the approximately expected time-average reward. The parameters for UCB1-LT and $(\alpha,\psi)$-UCB are respectively $u_{0}=1$ and $\alpha=2.5$ or $4$ with $\zeta=3600$ as their common parameter.
These results reveal an interesting phenomenon: the actual performance of UCB-LT is similar to that of $(\alpha,\psi)$-UCB which has a better theoretical regret bound over long-run for $\alpha=4$. However, a smaller choice of $\alpha$ slightly improves the performance of $(\alpha,\psi)$-UCB over long-run as consistent with the theoretical bound in (53) whose leading constant of the logarithmic order decreases as $\alpha$ decreases.
[Uncaptioned image][Uncaptioned image]
Figure 2: Performance comparison between UCB1-LT and $(\alpha,\psi)$-UCB
4 Conclusions
In this paper, We have proposed two UCB policies, namely the extended robust UCB and UCB1-LT, dealing with the heavy-tailed and light-tailed reward distributions in the frequentist multi-armed bandit problems, respectively.
The extended robust UCB policy aims at the class of heavy-tailed arm distributions and performs well in achieving the optimal logarithmic regret growth order. The policy takes advantage of a controlled relationship between different moments and discards the prerequisite knowledge of an upper bound of the moments. It is a generalization of the scale free algorithm proposed by Lattimore [21] with the assumption that $p=4$. In scenarios characterized by low discrimination and when $p=4$, our extended robust UCB policy exhibits a tighter regret upper bound in the long run. The achieved regret is merely a constant factor plus a constant term away from the established lower bound.
The UCB1-LT is efficient for the class light-tailed reward distributions where the optimal logarithmic regret growth order is attained. Through numerical evaluations, we observe a comparable actual performance to its extension, $(\alpha,\psi)$-UCB. Since the upper confidence bound in UCB1-LT is easy to compute (with $O(1)$ time complexity), it provides a practical archetype of UCB policies for light-tailed reward distributions.

### v4

@@
-thanks: (No relevant financial or non-financial competing interests. Derived from preprint [23].)
+
 Abstract
@@
 [5] and Lattimore [22].
-The previous UCB policies require the knowledge of an upper bound on specific moments of reward distributions or a particular moment to exist, which can be hard to acquire or guarantee in practical scenarios.
-Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p$ and $q$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of the UCB policies for the heavy-tailed reward distributions.
-keywords:
-frequentist multi-armed bandit problems; the light-tailed reward distributions; the heavy-tailed reward distributions; upper confidence bounds
+The previous UCB policies require some strict conditions on the reward distributions, which can be hard to guarantee in practical scenarios.
+Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p>q>1$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of the UCB policies for the heavy-tailed reward distributions.
+Furthermore, we achieve a near-optimal regret order without any knowledge of the reward distributions as long as their $p$-th moments exist for some $p>1$.
 1 Introduction
@@
 choosing an arm appearing to be best based on past observations.
-Under resource constraints, the way we acquire knowledge becomes crucial for the efficiency in maximizing the long-term reward in modern AI systems [31].
-While knowledge can be generally treated as data (usually not free in a strict sense), the struggle between exploration (sampling) and exploitation (consuming) as the prototypical dilemma in reinforcement learning is expressly reflected in an MAB model [28, 20, 3, 24, 5].
-Specifically, the original stochastic model of the multi-armed bandit problem considers a bandit machine with several arms; each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or, agent).
+Under resource constraints, the way we acquire knowledge becomes crucial in maximizing the long-term reward rate in modern AI systems [31].
+Since knowledge can generally be treated as data, the struggle between exploration (sampling) and exploitation (consuming) as the prototypical dilemma in reinforcement learning is expressly reflected in an MAB model [28, 21, 3, 24, 5].
+Specifically, the original stochastic model of the multi-armed bandit problem considers a bandit machine with several arms; each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or agent).
 Each reward distribution has a mean, say $\theta_{i}$.
 The player who has limited knowledge about the underlying reward distributions (no knowledge about the means $\{\theta_{i}\}$), has to choose only one of these arms to play and accrue a reward at each time step, given the reward history so far observed.
-The target of the player is to maximize the rate of gaining reward as the time horizon $T$ grows.
-In many applications of online decision-making problems, such as Internet advertising and recommendation systems, the MAB model is widely assumed with numerous algorithms proposed for efficient reinforcement learning, e.g., Zhao [38], Lattimore and Szepesvári [22].
-In a recommendation system, the vendor desires to know which item an online customer is looking for with the most interest.
-When we suggest a product and the customer makes a purchase, we receive a positive reward.
-Conversely, if the customer decides not to buy the recommended item, our reward may be either zero or negative, reflecting a potential waste of resources.
-Therefore the target of a recommendation system is to guess which items are most wanted by each customer, and the problem can be translated into the MAB model directly.
-To address the key issue in the conflict between exploration and exploitation, we note that it is impossible for the player to exactly know the expected values of the reward distributions within any finite horizon of time.
+The aim of the player is to maximize the reward rate that accrues as the time horizon $T$ grows.
+In many applications of online decision-making problems, such as Internet advertising and recommendation systems, the MAB model is widely assumed, with numerous algorithms proposed for efficient reinforcement learning, e.g.
+Zhao [38], Lattimore and Szepesvári [23].
+To address the key issue in the conflict between exploration and exploitation, we note that it is impossible for the player to exactly know the values of reward means within any finite horizon of time.
 The only information the player possesses is the (local) reward history, and the player has to learn from this to make a decision at each time step.
 Based on the limited historical information, the player has to choose between two options.
-The first option is to explore the arms with a lower average reward, as there is always a possibility that the best arm may not give the best average reward in the past sample path.
+The first option is to explore a suboptimal arm with a lower average reward, as there is always a possibility that the best arm may not give the best average reward in the past sample path.
 The second option is to exploit the arm with the highest average reward offered so far.
 Robbins [28] considered a simple two-armed bandit problem and showed that there is a policy to achieve the best average reward asymptotically as time goes to infinity.
-Many years later, Lai and Robbins [20] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the benchmark case that the player already knows which arm has the highest reward mean from the very beginning.
+Many years later,Lai and Robbins [21] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the benchmark scenario in which the player already knows which arm has the highest reward mean from the very beginning.
 Given a set of $K$ arms with reward distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$ with means $\{\theta_{i}\}_{i=1}^{K}$ and a policy $\pi$, the regret ~$\tilde{\mathcal{R}}_{\pi}^{\mathcal{F}}(T)$ is defined as (without loss of generality, we can assume that $\theta_{1}\geq\theta_{i}$ for $1\leq i\leq K$)
-$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=\sum_{\theta_{i}<\theta_{1}}E[s_{i}(T)]\Delta_{i},$ (1)
+$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=\sum_{\theta_{i}<\theta_{1}}\mathbb{E}[s_{i}(T)]\Delta_{i},$ (1)
 where
@@
 Furthermore, any sub-linear regret order implies the best expected time-average reward as $T\rightarrow\infty$.
-Strikingly, Lai and Robbins [20] proved an asymptotic lower bound on regret, with the growth rate of $O(\log T)$, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown.
+Strikingly, Lai and Robbins [21] proved an asymptotic $O(\log T)$ lower bound on the regret growth rate, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown.
 They also proposed an asymptotically optimal policy for a family of distributions, under which the best regret can be achieved asymptotically, both the logarithmic order and the leading coefficient.
-Later Burnetas and Katehakis [6] proved a generalized result under non-parametric settings.
-Suppose $\mathcal{H}$ is a set of distributions on $\mathbb{R}$.
-We call a bandit $\mathcal{F}$ (considered as a set of reward distributions of all arms) non-trivial, if there is at least one distribution (re-index it to $1$ if necessary) $F_{1}\in\mathcal{F}$ satisfying $\theta_{1}>\theta_{i}$ for all $i\in\{2,\dots,K\}$.
-They proved that for any policy $\pi$ s.t.
-for each fixed nontrivial $\mathcal{F}\subset\mathcal{H}$ (this policy condition is called consistency over $\mathcal{H}$)
-$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=o(T^{a}),\ \forall a>0,$ (3)
-we have that
-$\lim\inf\frac{\mathbb{E}[s_{i}(T)]}{\log T}\geq\sup\left\{\frac{1}{\text{KL}(F_{i},Y)}:Y\in\mathcal{H}\ \text{and}\ \mathbb{E}Y>\theta_{i}\right\},$ (4)
-where $\text{KL}(\cdot,\cdot)$ denotes the Kullback-Leibler distance between two distributions.
-This theorem demonstrates that for any consistent strategy under a given family of distributions, the regret will not be lower than the logarithmic growth.
 Auer et al.
-[3] proposed a class of policies named and featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $[a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player.
+[3] proposed a class of policies featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $[a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player.
 Their UCB1 and UCB1-Tuned polices perform very well for the classical MAB problem with a theoretically proven regret bound for any finite time horizon.
-The UCB policies are widely applied in machine learning development, e.g., the recommendations systems [32, 16], and in various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [19], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [15].
+UCB policies are widely applied in machine learning development (e.g.
+recommendation systems [32, 17]) and various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [20], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [16].
 1.1 Heavy-Tailed MAB Model
 The class of heavy-tailed distributions is a concept in contrast to the light-tailed case.
-Light-tailed (or, sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation, while heavy-tailed distributions have a relatively high probability of producing a large deviation.
+Light-tailed (or sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation, while heavy-tailed distributions have a relatively high probability of producing a large deviation.
 There are several popular light-tailed probability distributions, such as Bernoulli, Gaussian, Laplacian, and Exponential.
-Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions [10].
+Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions Chareka [10].
 Formally, a random variable $X$ is called light-tailed if there exists some $u_{0}>0$ such that its moment-generating function is well-defined on $[-u_{0},u_{0}]$ [10]:
 $M(u):=\mathbb{E}[\exp(uX)]<\infty,\ \ \forall|u|\leq u_{0}.$ (3)
-The above condition is equivalent to [10]
+The above condition is equivalent to [10]:
 $\mathbb{E}[\exp(u(X-\mathbb{E}X))]\leq\exp(\zeta u^{2}/2),\ \ \forall|u|\leq u_{0},\ \forall\zeta\geq\sup_{|u|\leq u_{0}}{M^{(2)}(u)},$ (4)
 where $M^{(2)}(\cdot)$ denotes the second derivative of $M(\cdot)$.
-Note that the above upper bound on $M(u)$ has a form of the moment generating function of the Gaussian distribution, thus the name “sub-Gaussian” has been adopted.
-The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution might take large values with higher probabilities, thus its moment-generating function does not exist (even locally).
-Compared to the light-tailed distributions, the heavy-tailed are harder to learn in terms of the rank of their means.
+Note that the above upper bound on $M(u)$ has a form of the moment-generating function of the Gaussian distribution, thus the name “sub-Gaussian” has been adopted.
+The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution can produce large values with high probabilities, leading to the nonexistence of its moment-generating function (even locally).
+Compared to the light-tailed class, the heavy-tailed one is harder to learn in terms of the rank of their means.
 Bubeck et al.
@@
 Robust UCB shows significant progress for the heavy-tailed MAB model by achieving the optimal logarithmic order of regret growth.
-Later on, Lattimore [21] developed a new UCB policy, which is referred to as the scale free algorithm.
+Subsequently, Lattimore [22] developed a new UCB policy referred to as the scale free algorithm.
 This algorithm removes the general assumption of a known upper bound on the $p$-th moment.
 Instead, it allows the variance and the squared fourth moment to scale freely as long as their ratio (kurtosis) is bounded by a known parameter.
-But this assumption restricts $p$ to be no less than $4$ (thus cannot be too heavy-tailed).
-In this paper, we extend Lattimore’s work to any $p>1$ with a tighter regret upper bound in the case of $p=4$ and low discrimination where arm means are close and learning the arm rank becomes challenging.
-From this point on, we assume that the player knows a constant $C_{p,q}$ for some $p$ and $q$ ($p>q>1$) such that
+But this assumption restricts $p$ to be no less than $4$ (not too heavy-tailed).
+In this paper, we extend Lattimore’s work to any $p>1$ with a tighter regret upper bound in the case of $p=4$ with low discrimination where arm means are close and learning the arm rank becomes challenging.
+First, we assume that the player knows a constant $C_{p,q}$ for some $p$ and $q$ ($p>q>1$) such that [^1] By Jensen’s inequality, the existence of the $p$-th moment implies the existence of the $q$-th moment for any $1<q<p$.
 $\mathbb{E}[|X-\mathbb{E}X|^{p}]\leq C_{p,q}\left(\mathbb{E}[|X-\mathbb{E}X|^{q}]\right)^{p/q}.$ (5)
-The above inequality is a direct generalization of the assumption on kurtosis in Lattimore [21] where $p=4$ and $q=2$.
-For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function [2]
+The requirement of knowledge of $C_{p,q}$ can be removed with an arbitrarily small sacrifice of the regret order as seen in Sec.
+2.1.
+The above inequality is a direct generalization of the assumption of kurtosis in Lattimore [22] where $p=4$ and $q=2$.
+For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function Arnold [2]
 $F(x)=\begin{cases}1-\left(1+\left(\frac{x-a}{\sigma}\right)^{1/\gamma}\right)^{-1},&x>a\\ 0,&x\leq a\end{cases},$
@@
 $C(p,q,\gamma)$ $=I(p,\gamma)/I(q,\gamma)^{p/q}.$
-However, the moment under this distribution is a function of $\sigma$; when $\sigma$ increases, the moment of any fixed order also increases, if it exists.
-The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment [5].
+However, the moment of this distribution is a function of $\sigma$.
+When $\sigma$ increases, the moment of any fixed order also increases, if it exists.
+The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment Bubeck et al.
+[5].
 Therefore, for an arm with an unknown $\sigma$, we cannot apply the robust UCB policy.
-Furthermore, the learning efficiency can be improved if a larger $p$ can be used (tighter bound on the tail, but the corresponding moment bound becomes harder to know or compute a priori), it is desired to eliminate the need for any prior knowledge regarding the specific values or bounds of the moments.
+Furthermore, learning efficiency can be improved if a larger $p$ can be used (tighter bound on the tail, but the corresponding moment bound becomes harder to know or compute a priori).
+It is thus better to eliminate the need for any prior knowledge regarding the specific values or bounds on the moments.
 This motivates Lattimore [22] to consider a scale free algorithm.
-After we generalize Lattimore’s work, we will give a correction to a minor flaw in a lemma of Bubeck et al.
-[5] which subsequently caused a small error in the computation of regret upper bound in Lattimore [21].
-1.2 Related Works
-The theorems of the frequentist MAB have been developed for a long time and various milestones have been made since Lai and Robbins [20].
+After we generalize Lattimore’s work, we will give a correction to a minor flaw of a lemma in Bubeck et al.
+[5] that subsequently caused a small error in the computation of the regret upper bound in Lattimore [22].
+1.2 Related Work
+The theorems of the frequentist MAB have been developed for a long period of time and various milestones have been made since Lai and Robbins [21].
 In contrast to the Bayesian model, the frequentist MAB does not assume a priori knowledge of the initial probability or state of the system but learns the core parameters solely through the past sampling history, e.g., computing an upper confidence bound as a function of the frequency that an arm is selected and the values of the observed samples.
-Such a way of decision-making in balancing exploration with exploitation contrasts to the Bayesian MAB where the tradeoff is usually addressed by dynamic programming for the time-varying characteristics of the system or the observation model [12].
+This way of decision-making in balancing exploration with exploitation contrasts with the Bayesian model in which the tradeoff is usually addressed by dynamic programming for the transitional characteristics of the system or the observation model [13].
 After the UCB1 policy by Auer et al.
-in [3], many other policies of the UCB-type have been proposed in the literature.
+[3], many other policies of the UCB class have been proposed in the literature.
 Maillard et al.
 [26] directed their attention to the prospect of KL divergence between probability distributions.
-They introduced the KL-UCB algorithm, which notably enhanced the regret bound.
-This improvement was achieved under the assumption of rewards being bounded and having a known upper bound.
+They introduced the KL-UCB algorithm, which notably tightened the regret bound.
+This improvement was achieved under the assumption of bounded rewards with a known bound.
 Furthermore, KL-UCB policy achieves the asymptotic lower bound of regret growth for this class of rewards given in Burnetas and Katehakis [6].
@@
 These policies were neither implemented nor proved to achieve the logarithmic regret growth in the case of unbounded reward distributions.
-In 2011, the first author of this paper proposed UCB1-LT with a proven bound on the regret growth of the logarithmic order for the class of light-tailed reward distributions [24], filling the vacancy for the case of unbounded light-tailed reward distributions.
+In 2011, the first author of this paper proposed UCB1-LT with a proven bound on the regret growth of the logarithmic order for the class of light-tailed reward distributions Liu and Zhao [24], filling the gap for the case of unbounded light-tailed reward distributions.
 Together with the extended robust UCB for the heavy-tailed class focused in this paper, we provide a relatively complete picture of UCB policies for general reward distributions.
-After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
+After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the reward moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
 For the case of heavy-tailed reward distributions, we have mentioned that Bubeck et al.
@@
 The robust UCB policy uses different mean estimators rather than the empirical one and remarkably achieves the logarithmic regret growth rate.
-Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator [8].
+Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator Catoni [8].
 The truncated mean estimator requires the knowledge of an upper bound of origin moments, which is not very effective in the case that the reward mean is far away from zero.
-The median-of-means estimator considers the (central) moment and thus patches up the deficiency of the truncated mean estimator: the mean estimation will not be affected when the reward distributions are changed by translations.
-The Catoni mean estimator is more complicated and can only be used when variances of the reward distributions exist ($p\geq 2$).
+The median-of-means estimator considers the (central) moment and thus patches up the deficiency of the truncated mean estimator: the mean estimation will not be affected when the reward distributions are transformed by translations.
+The Catoni mean estimator is more complicated and can only be used when the reward distributions have variances ($p\geq 2$).
 But when applied in the robust UCB policy, it gives a much smaller regret coefficient than the other two estimators.
-Recently Chen et al.
-[11] showed a result of applying Catoni mean estimator to the case that $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators.
+Recently, Chen et al.
+[11] showed a result of applying Catoni mean estimator to the case of $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators.
 Finer estimations when extending the Catoni mean estimator to the case of $p<2$ are interesting for future investigations.
@@
 Agrawal et al.
-[1] established a policy called $\text{KL}_{\inf}$-UCB, which achieves a logarithmic upper bound of regret very close to the asymptotic lower bound proved by Burnetas and Katehakis [6].
+[1] established a policy called $\text{KL}_{\inf}$-UCB, which achieves a logarithmic upper bound on regret very close to the asymptotic lower bound proved by  Burnetas and Katehakis [6].
 For other non-classical MAB models, the heavy-tailed model has also been considered [37, 30].
-However, all these policies do not discard the knowledge requirement of an upper bound of moments, which has been removed from the extended robust UCB policy proposed in this paper.
+However, these policies do not eliminate the knowledge requirement of an upper bound of moments, which has been removed from the extended robust UCB policy proposed in this paper.
+Note that Genalti et al.
+[12] considered the case where no knowledge of the moments is available except for the existence of moments of order greater than $1$, but a relatively strong assumption of the distribution (a larger probability mass on the negative semi-axis) was needed to achieve logarithmic regret.
 There are also policies not following the idea of UCB.
-The $\varepsilon$-greedy policy is well-studied, and many policies are proposed using its idea, e.g., the constant $\varepsilon$-decreasing policy [34], GreedyMix [9] and $\varepsilon_{n}$-greedy [3].
-The Deterministic Sequencing of Exploration and Exploitation (DSEE) policy [25, 33] achieves the logarithmic regret for both cases of single player and multiple players (with the knowledge on moment upper bounds).
-Other recent progresses on variants of MAB, such as adversarial bandits, contextual bandits and linear bandits, can be found in Lattimore and Szepesvári [22, 13, 17].
-2 The Heavy-Tailed Distribution
+The $\epsilon$-greedy policy was well studied, and many policies were proposed using its idea, e.g., the constant $\epsilon$-decreasing policy [34], GreedyMix [9], and $\epsilon_{n}$-greedy [3].
+The Deterministic Sequencing of Exploration and Exploitation (DSEE) policy [25, 33] also achieves logarithmic regret for both cases of single player and multiple players (with knowledge of upper bound on moments).
+Other recent progresses on variants of MAB, such as adversarial bandits, contextual bandits, and linear bandits, can be found in [23, 14, 18].
+2 Heavy-Tailed Reward Distributions
 2.1 The Extended Robust UCB Policy
-Consier $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$.
-In contrast with the light-tailed class, the heavy-tailed distribution may yield a very high (or low) reward realization with a high probability.
+Consider $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with heavy-tailed distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$.
+In contrast with the light-tailed class, a heavy-tailed distribution may yield a very high (or low) reward realization with a high probability.
 Consequently, the empirical mean estimator $\overline{X}_{i,s}=\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for a sequence of i.i.d.
-random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work well as an estimation to the expectation of $X_{i}$ for arm $i$.
+random variables $\{X_{i,j}\}_{j=1}^{s}$ may not work properly to estimate the expectation of $X_{i}$ for arm $i$.
 The appendix of  Bubeck et al.
@@
 $\frac{1}{s}\sum_{j=1}^{s}X_{i,j}\leq\theta_{i}+\left(\frac{3\mathbb{E}|X_{i}-\theta_{i}|^{r}}{\delta s^{r-1}}\right)^{\frac{1}{r}},$
-where $1<r\leq 2$.
-This bound is too loose to obtain a UCB policy achieving the logarithmic regret growth.
-To remedy this issue of divergence, we adopt the median-of-means estimator proposed in Bubeck et al.
-[5].
+where $1<r\leq 2$. This bound is too loose to obtain a UCB policy for achieving logarithmic regret growth. To remedy this issue, we adopt the median-of-means estimator proposed in  Bubeck et al. [5].
 Definition 1.
@@
 Lemma 1.
-Given an arbitrary $\varepsilon>0$. Suppose that
-$\log\delta^{-1}>1/\varepsilon,\ s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\varepsilon\log(\delta^{-1})-1}$
-and $v=\mathbb{E}[|X_{i}-\theta_{i}|^{p}]<\infty$, for some $p$ such that $p>1$. Then with probability at least $1-\delta$, the median-of-means estimator
-$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\geq\theta_{i}-(12v)^{1/p}\left(\frac{(8+\varepsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
-Also, with a probability of at least $1-\delta$,
-$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\leq\theta_{i}+(12v)^{1/p}\left(\frac{(8+\varepsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
-Proof.
-Observe that $s\geq\lceil 8\log\delta^{-1}\rceil$ holds under the given condition in the lemma. For simplicity of presentation, we introduce some symbols as below. Let
-$k$ $:=\lceil 8\log\delta^{-1}\rceil,$ (8)
-$\chi$ $:=\frac{(8+\varepsilon)\log\delta^{-1}}{s},$ (9)
-$\eta$ $:=(12v)^{1/p}\chi^{(p-1)/p},$ (10)
-$\hat{\mu}_{l}$ $:=\frac{1}{N}\sum_{j=lN+1}^{(l+1)N}X_{i,j},\ \ l=0,1,\dots,k-1.$ (11)
-According to the appendix of Bubeck et al. [5],
-$\xi:=\mathbb{P}\left(\hat{\mu}_{l}>\theta_{i}+\eta\right)\leq\frac{3v}{N^{p-1}\eta^{p}}=\frac{1}{4N^{p-1}\chi^{p-1}}.$
-Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$.
-Since $s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\varepsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$.
-Then using Hoeffding’s inequality for the tail of a binomial distribution [14], we have
-$\mathbb{P}\left(\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)>\theta_{i}+\eta\right)\leq\mathbb{P}\left(\sum_{l=0}^{k-1}\mathbb{I}(\hat{\mu}_{l}\geq\theta_{i}+\eta)\geq\frac{k}{2}\right)\leq\exp(-2k(1/2-\xi)^{2})\leq\delta.$
-Similarly, the other inequality also holds by symmetry. ∎
+Given any $\epsilon>0$. Suppose that
+$\log\delta^{-1}>1/\epsilon,\ s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$
+and $v=\mathbb{E}[|X_{i}-\theta_{i}|^{p}]<\infty$ for some $p$ such that $p>1$. Then with probability at least $1-\delta$, the median-of-means estimator
+$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\geq\theta_{i}-(12v)^{1/p}\left(\frac{(8+\epsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
+Also, with probability at least $1-\delta$,
+$\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},\lceil 8\log\delta^{-1}\rceil)\leq\theta_{i}+(12v)^{1/p}\left(\frac{(8+\epsilon)\log(\delta^{-1})}{s}\right)^{(p-1)/p}.$
 With the help of the median-of-means estimator, we can now define the upper confidence bound for the extended robust UCB policy.
-As mentioned above, we assume that for any arm distribution $X_{i}\in\mathcal{F}$ with mean $\theta_{i}$, there is a moment control coefficient $C_{p,q}$ ($1<q<p$) such that
+As mentioned in (5), we first assume that for any arm distribution $X_{i}\in\mathcal{F}$ with mean $\theta_{i}$, there is a known moment control coefficient $C_{p,q}$ ($1<q<p$) such that
 $\mathbb{E}[|X_{i}-\theta_{i}|^{p}]\leq C_{p,q}\mathbb{E}[|X_{i}-\theta_{i}|^{q}]^{r},$ (6)
@@
 Definition 2.
-Recall the notations introduced by (8) and (9).
-Given an arbitrary $\varepsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
+Let $k=\lceil 8\log\delta^{-1}\rceil,\chi=\frac{(8+\epsilon)\log\delta^{-1}}{s}$.
+Given an arbitrary $\epsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
 $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta):=$
@@
 $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)\geq\theta_{i}.$
-Proof.
+Corollary 1.
+Under the conditions of Lemma 1, with probability at least $1-\delta$, we have
+$\hat{\mu}(\{Y_{i,j}\}_{j=1}^{s},k)-v_{q}\leq C^{\prime}v_{q}\chi^{(p-q)/p}.$
+Please see (27) for the definitions of $Y_{i,j},v_{q},Y_{i},v_{p}$.
+Finally, with the upper confidence bound $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s})$, we complete the extended robust UCB policy as follows.
+Input: $\epsilon>0$, $K$: the number of arms
+Initialize: $s_{i}\leftarrow 0$ and $\{X_{i,j}\}_{j=1}^{s_{i}}$ an empty sequence for each arm $i$, $1\leq i\leq K$
+for $t\leftarrow 1$ to $T$ do
+       for $i\leftarrow 1$ to $K$ do
+             if $s_{i}=0$ then
+                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
+             end if
+            else
+                  Compute arm $i$’s upper confidence bound as $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$
+             end if
+       end for
+      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
+       Update: $s_{j}\leftarrow s_{j}+1$ and append $x$ to sequence $\left\{X_{j,k}\right\}_{k=1}^{s_{j}}$
+end for
+Algorithm 1 The Extended Robust UCB Policy $\pi^{*}$
+Here we present the main theorem that the extended robust UCB policy achieves the logarithmic order of regret growth.
+Theorem 1.
+Suppose $p>q>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$ and (6) is satisfied with known $C_{p,q}$.
+Then for any $\epsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
+To prove Theorem 1, we just need to show the following core lemma to bound the time $s_{i}$ spent on each suboptimal arm $i$ by the extended robust UCB policy.
+Lemma 3.
+The expected number of times $\mathbb{E}[s_{i}(T)]$ that a suboptimal arm is chosen by the extended robust UCB policy $\pi^{*}$ has an upper bound of the logarithmic order with $T$.
+A Monte Carlo simulation example based on the Pareto distribution of type III (see Sec.
+1.1) is shown in Fig.
+1.
+Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e.
+$\chi$ defined in (24) becomes sufficiently small such that, within the $\sup$ operator of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is not zero and the growth rate of the right-hand side with $\theta$ becomes small).
+Note that if $C_{p,q}$ is unknown, then the player can simply choose any increasing function $f$ with $f(t)\rightarrow+\infty$ as $t\rightarrow+\infty$ to replace $C_{p,q}$ and achieve a regret arbitrarily close to logarithmic order without requiring any knowledge of the reward moments (see Appendix F for a proof).
+Figure 1: The performance of extended robust UCB
+2.2 Further Analysis and Comparison of The Extended Robust UCB
+Now we consider scenarios of specific chosen $p$ and $q$ to further improve the efficiency of the extended robust UCB policy.
+Case 1.
+Suppose that the $p^{\prime}$-th order of moments of reward distributions exists with $p^{\prime}>2$ for all arms.
+Then we can choose $p$ and $q$ such that $p=2q\leq p^{\prime}$.
+The inequality (29) can be refined to
+$\begin{split}\mathbb{E}[|Y_{i}-v_{q}|^{p/q}]=\mathbb{E}Y_{i}^{2}+v_{q}^{2}-2v_{q}\mathbb{E}Y_{i}=\mathbb{E}Y_{i}^{2}-v_{q}^{2}=v_{p}-v_{q}^{2}\leq(C_{p,q}-1)v_{q}^{p/q}.\end{split}$ (8)
+Hence, we can refine $C^{\prime}$ from (7) to
+$C^{\prime}=\sqrt{12(C_{p,q}-1)},$ (9)
+which can be much smaller than the original definition (7).
+Now we look at the definition of $\chi_{M}$ in Appendix E that ensures the three inequalities (39), (40) and (41) to hold.
+The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e.
+the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy’s selection will converge to the best arm faster.
+Case 2. Assume that the $4$-th order of moments of reward distributions exists as in  Lattimore [22]. Then we can choose $p=4$ and $q=2$. In this case, our UCB in Definition 2 can be rewritten as
+[…]
+In other words, $C_{p,q}$ degenerates to an upper bound on kurtosis. The function $B(x)$ defined in (36) becomes
+$B(x)=\frac{24\tau(x)}{1-48x\tau(x)}\left(1+C^{\prime}\sqrt{x}\right)+\frac{576\sqrt{x}\tau(x)}{1-48x\tau(x)},$
+and the three inequalities (39), (40) and (41) are refined to
+11 11 $>C^{\prime}\sqrt{x},$ (10)
+11 11 $>48x\tau(x),$ (11)
+$\Delta_{i}$ $\geq\sqrt{v_{2}x}\left(\sqrt{B(x)}+2\sqrt{3}\right).$ (12)
+Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (46) is significantly decreased.
+Now we compare the extended robust UCB and the scale free algorithm proposed in  Lattimore [22], under the same assumption specified at the beginning of Case 2.
+We first address a flaw of Lemma 2 in Bubeck et al.
+[5].
+This lemma, crucial in Lattimore [22] for deriving the regret upper bound of the scale free algorithm, is examined more closely below.
+For clarity and coherence with the notation employed in this paper, we present the following adapted version.
+A Flawed Statement  Bubeck et al.
+[5]: Let $\delta\in(0,1)$ and $p\in(1,2]$.
+Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d.
+random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $v_{p}$.
+Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$.
+Then with probability at least $1-\delta$,
+$\hat{\mu}\left(\{X_{i,j}\}_{j=1}^{s},k^{\prime}\right)\leq\theta_{i}+(12v_{p})^{1/p}\left(\frac{16\log({\rm e}^{1/8}\delta^{-1})}{s}\right)^{(p-1)/p}.$
+The last inequality in the proof of this statement in  Bubeck et al.
+[5] assumes that $\exp(-k^{\prime}/8)\leq\delta$, equivalent to $k^{\prime}\geq 8\log\delta^{-1}$.
+However, the definition of $k^{\prime}$ in this statement is
+$k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor.$
+In the following, we give the corrected version of the above statement.
+Lemma 4.
+Let $\delta\in(0,1)$ and $p\in(1,2]$.
+Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d.
+random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $v_{p}$, where $s$ is chosen such that $\lfloor s/2\rfloor\geq 8\log\delta^{-1}$.
+Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$.
+Then with probability at least $1-\delta$,
+$\hat{\mu}\left(\{X_{i,j}\}_{j=1}^{s},k^{\prime}\right)\leq\theta_{i}+(12v_{p})^{1/p}\left(\frac{16\log({\rm e}^{1/8}\delta^{-1})}{s}\right)^{(p-1)/p}.$
+The next theorem shows that with a properly chosen parameter $\epsilon>0$, the upper bound given in (46) is also tighter than that in Lattimore [22] in long run if $\Delta_{i}$ is small, i.e., arms are hard to be distinguished (the low-discrimination case).
+In the following theorem, we formally state this result while correcting the upper bound in Lattimore [22] by Lemma 4.
+Theorem 2.
+Choose any $8<\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$ ($\approx 44.158$) and let $\delta=t^{-2}$.
+The upper bound given in Lattimore [22] of $\mathbb{E}[s_{i}(T)]$ is greater than that given in (46) for sufficiently large $T$ and sufficiently small $\Delta_{i}$.
+That is, after dividing both regret upper bounds by $\log\delta^{-1}$ and letting $T\to+\infty$, we have
+$(8+\epsilon)\left(\frac{8}{\epsilon}+\max\left\{\frac{8}{\epsilon},\frac{1}{\chi_{M}}\right\}\right)<3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}+16$ (13)
+for sufficiently small $\Delta_{i}$.
+The regret upper bound of the scale free algorithm proposed in  Lattimore [22] is thus asymptotically larger than that of the extended robust UCB policy for $\epsilon\in(8,\frac{280-16\sqrt{2}}{2\sqrt{2}+3})$.
+Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound (in the non-parametric setting) by only a constant factor and a constant term as in the scale free algorithm of  Lattimore [22].
+First, we restate the lower bound derived in  Lattimore [22] as follows.
+Theorem 3.
+Let $H_{\kappa_{0}}$ be the set of distributions that have kurtosis less than $\kappa_{0}$.
+Assume $\Delta>0$ and $\kappa_{0}\geq 7/2$.
+Suppose that $X\in H_{\kappa_{0}}$ has a mean of $\mu$, a positive variance of $\sigma^{2}$, and a kurtosis of $k$.
+Then
+$\begin{split}\inf\{{\rm KL}(X,X^{\prime}):&X^{\prime}\in H_{\kappa_{0}}\ \text{and}\ \mathbb{E}[X^{\prime}]>\mu+\Delta\}\\ \leq&\begin{cases}\min\{-\log(1-p),\frac{C_{1}\Delta^{2}}{\sigma^{2}}\},&\text{if}\ C_{0}\sqrt{k}(k+1)\frac{\Delta}{\sigma}<\kappa_{0}\\ -\log(1-p),&\text{otherwise}\end{cases},\end{split}$ (14)
+[…]
+Based on this, we can draw the following conclusion:
+Theorem 4.
+Let $\delta=t^{-2}$.
+Under Case 2, the regret of the extended robust UCB policy applied to $H_{\kappa_{0}}$ ($\kappa_{0}\geq 7/2$) differs from the lower bound of regret by a constant factor and a constant term.
+3 Light-Tailed Reward Distributions
+We point out that our policy and results also work for the light-tailed distributions since our construction depends only on the existence of moments.
+However, the leading constant of the logarithmic regret order can be improved by our past work in Liu and Zhao [24], which we formally and briefly discuss here with a comparison with some subsequent work following Liu and Zhao [24].
+3.1 The UCB1-LT Policy
+The empirical mean is adopted in estimating the upper confidence bound for the class of light-tailed reward distributions.
+As before, let $\overline{X}_{i,s}$ denote the empirical mean $\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for arm $i$, where $\{X_{i,j}\}_{j=1}^{s}$ form the i.i.d.
+random reward sequence drawn from an unknown distribution $X_{i}\in\mathcal{F}$.
+Lemma 5.
+(Bernstein-type bound) For i.i.d.
+random variables $\{X_{i,j}\}_{j=1}^{s}$ drawn from a light-tailed distribution $X_{i}$ (with mean $\theta_{i}$) with a finite moment-generating function $M(u)$ over range $u\in[-u_{0},u_{0}]$ for some $u_{0}>0$.
+We have, $\forall\epsilon>0$,
+$\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\geq\epsilon)\leq\begin{cases}\exp(-\frac{s}{2\zeta}\epsilon^{2}),&\epsilon<\zeta u_{0}\\ \exp(-\frac{su_{0}}{2}\epsilon),&\epsilon\geq\zeta u_{0}\\ \end{cases},$ (15)
+where $\zeta>0$ satisfies $\zeta\geq\sup_{|u|\leq u_{0}}{M^{(2)}(u)}$. A similar bound also holds for $\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\leq-\epsilon)$ by symmetry.
+The proof of the above lemma follows a similar argument as in  Vershynin [35]. Using the Bernstein-type bound by Lemma 5, we propose the UCB1-LT policy as follows.
+Input: $a_{1}\geq 8\zeta$, $a_{2}\geq a_{1}/(\zeta u_{0})$, $K$: the number of arms
+Initialize: $s_{i}\leftarrow 0$ and $\overline{X}_{i}\leftarrow 0$ for each arm $i$, $1\leq i\leq K$
+for $t\leftarrow 1$ to $T$ do
+       for $i\leftarrow 1$ to $K$ do
+             if $s_{i}=0$ then
+                  Assign +$\infty$+ ∞ to the upper confidence bound
+             end if
+            else
+                   if $\sqrt{\frac{a_{1}\log t}{s_{i}}}<\zeta u_{0}$ then
+                         Compute the upper confidence bound as $\overline{X}_{i}+\sqrt{\frac{a_{1}\log t}{s_{i}}}$
+                   end if
+                  else
+                         Compute the upper confidence bound as $\overline{X}_{i}+\frac{a_{2}\log t}{s_{i}}$
+                   end if
+             end if
+       end for
+      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
+       $\overline{X}_{j}\leftarrow\frac{x+s_{j}\overline{X}_{j}}{s_{j}+1}$ and $s_{j}\leftarrow s_{j}+1$
+end for
+Algorithm 2 The UCB1-LT Policy
+The UCB1-LT policy considers two upper confidence bounds and alternatively uses one of them according to values of $t$ and $s_{i}(t-1)$.
+To minimize the theoretical regret upper bound, we can choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$.
+The following theorem shows that UCB1-LT achieves logarithmic regret growth for the light-tailed reward distributions.
+Theorem 5.
+For the light-tailed reward distributions, the regret of the UCB1-LT policy satisfies the following inequality:
+$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{i:\theta_{i}<\theta_{1}}\Delta_{i}\left(\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}\log T+1+\frac{\pi^{2}}{3}\right).$ (16)
+In Appendix A, we compare our UCB1-LT policy with the $(\alpha,\psi)$-UCB policy proposed in Bubeck and Cesa-Bianchi [4].
+4 Conclusion and Acknowledgment
+In this paper, we have proposed two order-optimal UCB policies, namely the extended robust UCB and UCB1-LT, dealing with the heavy-tailed and light-tailed reward distributions in the frequentist multi-armed bandit problems, respectively.
+We are grateful for the help from Mr.
+Haoran Chen during the initial stage of this project, and Dr.
+Yaoqing Yang from Dartmouth College in improving this article.
+Appendix A From UCB1-LT to $(\alpha,\psi)$-UCB
+After the establishment of UCB1-LT by Liu and Zhao [24],  Bubeck and Cesa-Bianchi [4] subsequently proposed the $(\alpha,\psi)$-UCB under a more general assumption on the moment generating functions.
+Specifically, the $(\alpha,\psi)$-UCB policy assumes that the distribution of reward $X_{i}\in\mathcal{F}$ satisfies the following condition: there exists a convex function $\psi$ defined on $\mathbb{R}$ such that for all $\lambda>0$,
+$\log\mathbb{E}\left[{\rm e}^{\lambda(X_{i}-\theta_{i})}\right]\leq\psi(\lambda)\text{ and }\log\mathbb{E}\left[{\rm e}^{\lambda(\theta_{i}-X_{i})}\right]\leq\psi(\lambda).$ (17)
+For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (17) becomes the well-known Hoeffding’s lemma [4].
+Since +$\infty$+ ∞ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT.
+The Legendre-Frenchel transform $\psi^{*}$ of $\psi$ is defined as
+$\psi^{*}(\epsilon):=\sup_{\lambda\in\mathbb{R}}\lambda\epsilon-\psi(\lambda).$
+The upper confidence bound of $(\alpha,\psi)$-UCB at time $t$ is defined as
+$I_{i}(t):=\overline{X}_{i,s_{i}(t-1)}+(\psi^{*})^{-1}\left(\frac{\alpha\log t}{s_{i}(t-1)}\right).$ (18)
+Theorem 6 (Bubeck and Cesa-Bianchi [4]).
+Assume that the reward distributions satisfy (17). Then $(\alpha,\psi)$-UCB with $\alpha>2$ achieves
+$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{\theta_{i}<\theta_{1}}\left(\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\log T+\frac{\alpha}{\alpha-2}\right).$ (19)
+For the UCB1-LT policy, we can choose
+$\psi(\lambda)=\begin{cases}\frac{\zeta\lambda^{2}}{2},&\lambda\leq u_{0}\\ +\infty,&\lambda>u_{0}\end{cases}.$ (20)
+Then (17) becomes the same assumption as in UCB1-LT (see (4)). In this case, we have
+$\psi^{*}(\epsilon)=\begin{cases}\frac{\epsilon^{2}}{2\zeta},&|\epsilon|<\zeta u_{0}\\ u_{0}(|\epsilon|-\frac{\zeta u_{0}}{2}),&|\epsilon|\geq\zeta u_{0}\end{cases}.$
+Thus the inverse $(\psi^{*})^{-1}(x)$ of $\psi^{*}$ for $x\geq 0$ will be
+$(\psi^{*})^{-1}(x)=\begin{cases}\sqrt{2\zeta x},&0\leq x<\frac{\zeta u_{0}^{2}}{2}\\ \frac{x}{u_{0}}+\frac{\zeta u_{0}}{2},&x\geq\frac{\zeta u_{0}^{2}}{2}\end{cases}.$
+Furthermore, if we choose $\alpha=4$, then the upper confidence bound defined in (18) becomes
+$I_{i}(t)=\overline{X}_{i,s_{i}(t-1)}+\begin{cases}\sqrt{\frac{8\zeta\log t}{s_{i}(t-1)}},&\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}\\ \frac{4\log t}{u_{0}s_{i}(t-1)}+\frac{\zeta u_{0}}{2},&\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}\end{cases}.$ (21)
+If we choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$ in UCB1-LT, we get the same UCB under the condition $\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}$.
+For the case $\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}$, the upper confidence bound of $(\alpha,\psi)$-UCB is not larger than that of UCB1-LT and thus may achieve a smaller theoretical upper bound of regret as follows:
+$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{\theta_{i}<\theta_{1}}\left(r(\Delta_{i})\log T+2\right),$ (22)
+where
+$r(\Delta_{i})=\begin{cases}\frac{32\zeta}{\Delta_{i}},&\Delta_{i}<2\zeta u_{0}\\ \frac{8\Delta_{i}}{u_{0}(\Delta_{i}-\zeta u_{0})},&\Delta_{i}\geq 2\zeta u_{0}\end{cases}.$
+Note that this regret bound is better than (16) in Theorem 5 for UCB1-LT as $T$ becomes sufficiently large, as shown in the following lemma.
+Lemma 6.
+Choose the parameter $\alpha$ in $(\alpha,\psi)$-UCB such that $\alpha\leq 4$, and assume that Assumption (4) is satisfied and thus $\psi(\lambda)$ is chosen as in (20).
+Then the logarithmic regret upper bound in Theorem 6 has a leading constant no larger than that in Theorem 5, i.e.
+$\sum_{\theta_{i}<\theta_{1}}\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\leq\sum_{\theta_{i}<\theta_{1}}\Delta_{i}\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}.$
+Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments.
+Assume that the arms are offering rewards according to $19$ normal distributions.
+The means of reward distributions are $0.5,1.0,1.5,\dots,9.5$ with standard deviations all set to $60$.
+A Monte Carlo (MC) simulation with 100 runs and time period $T=50000$ is shown below.
+The horizontal axis in both figures is the time step $t$.
+The first figure only shows $T$ from $10000$ to $50000$ for better display, while the second one ranges from $1$ up to $50000$.
+The vertical axis of the left figure is the logarithmic regret averaged from the 100 MC runs (to approximate the expectation), while of the right one is the approximately expected time-average reward.
+The parameters for UCB1-LT and $(\alpha,\psi)$-UCB are respectively $u_{0}=1$ and $\alpha=2.5$ or $4$ with $\zeta=3600$ as their common parameter.
+These results reveal an interesting phenomenon: the actual performance of UCB-LT is similar to that of $(\alpha,\psi)$-UCB which has a better theoretical regret bound over long-run for $\alpha=4$.
+However, a smaller choice of $\alpha$ slightly improves the performance of $(\alpha,\psi)$-UCB over long-run as consistent with the theoretical bound in (19) whose leading constant of the logarithmic order decreases as $\alpha$ decreases.
+Figure 2: Performance comparison between UCB1-LT and $(\alpha,\psi)$-UCB
+Appendix B Proof of Lemma 1
+Observe that $s\geq\lceil 8\log\delta^{-1}\rceil$ holds under the given condition in the lemma. For simplicity of presentation, we introduce some symbols as below. Let
+$k$ $:=\lceil 8\log\delta^{-1}\rceil,$ (23)
+$\chi$ $:=\frac{(8+\epsilon)\log\delta^{-1}}{s},$ (24)
+$\eta$ $:=(12v)^{1/p}\chi^{(p-1)/p},$ (25)
+$\hat{\mu}_{l}$ $:=\frac{1}{N}\sum_{j=lN+1}^{(l+1)N}X_{i,j},\ \ l=0,1,\dots,k-1.$ (26)
+According to the appendix of  Bubeck et al. [5],
+$\xi:=\mathbb{P}\left(\hat{\mu}_{l}>\theta_{i}+\eta\right)\leq\frac{3v}{N^{p-1}\eta^{p}}=\frac{1}{4N^{p-1}\chi^{p-1}}.$
+Note that $N=\lfloor s/\lceil 8\log\delta^{-1}\rceil\rfloor$.
+Since $s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$.
+Then using Hoeffding’s inequality for the tail of a binomial distribution Hoeffding [15], we have
+$\mathbb{P}\left(\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)>\theta_{i}+\eta\right)\leq\mathbb{P}\left(\sum_{l=0}^{k-1}\mathbb{I}(\hat{\mu}_{l}\geq\theta_{i}+\eta)\geq\frac{k}{2}\right)\leq\exp(-2k(1/2-\xi)^{2})\leq\delta.$
+Similarly, the other inequality also holds by symmetry. $\hfill\qed$
+Appendix C Proof of Lemma 2
 Recall the notations introduced by (23), (24) and (7).
@@
 $\theta_{i}\leq\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)+\left(\frac{12\hat{\mu}(\{|X_{i,j}-\theta_{i}|^{q}\}_{j=1}^{s},k)}{\max\left\{0,1-C^{\prime}\chi^{(p-q)/p}\right\}}\right)^{1/q}\chi^{(q-1)/q}.$
-The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. ∎
-Corollary 1.
-Under the conditions of Lemma 1, with probability at least $1-\delta$, we have
-$\hat{\mu}(\{Y_{i,j}\}_{j=1}^{s},k)-v_{q}\leq C^{\prime}v_{q}\chi^{(p-q)/p}.$
-Proof.
-This corollary is a direct consequence of Lemma 1 and the inequality in (16). ∎
-Finally, with the upper confidence bound estimator $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s})$, we complete the extended robust UCB policy as follows.
-1 Input: $\varepsilon>0$, $K$: the number of arms
-2 Initialize: $s_{i}\leftarrow 0$ and $\{X_{i,j}\}_{j=1}^{s_{i}}$ an empty sequence for each arm $i$, $1\leq i\leq K$
-3 for $t\leftarrow 1$ to $T$ do
-4       for $i\leftarrow 1$ to $K$ do
-5             if $s_{i}=0$ then
-6                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
-7             end if
-8            else
-9                  Compute arm $i$’s upper confidence bound as $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$
-10             end if
-11
-12       end for
-13      Choose an arm $j$ that maximizes the upper confidence bound and obtain reward $x$
-14       Update: $s_{j}\leftarrow s_{j}+1$ and append $x$ to sequence $\left\{X_{j,k}\right\}_{k=1}^{s_{j}}$
-15 end for
-Algorithm 1 The Extended Robust UCB Policy
-Here we present the main theorem that the extended robust UCB policy achieves the logarithmic order of regret growth.
-Theorem 1.
-Suppose $p>1$ such that the $p$-th order moments exist for all $F_{i}\in\mathcal{F}$, $1<q<p$ and $C_{p,q}$ such that (12) holds.
-Then for any $\varepsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
-To prove Theorem 1, we establish the following core lemma to bound the time $s_{i}$ spent on each arm $i$ by the extended robust UCB policy.
-Lemma 3.
-The expected number of times $\mathbb{E}[s_{i}(T)]$ that a sub-optimal arm is chosen by the extended robust UCB has an upper bound of the logarithmic order with $T$.
-Proof.
+The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. $\hfill\qed$
+Appendix D Proof of Corollary  1
+This corollary is a direct consequence of Lemma 1 and the inequality in (29). $\hfill\qed$
+Appendix E Proof of Lemma 3
 In addition to notations (23), (24) and (7), we introduce two new notations:
@@
 $|a+b|^{q}\leq 2^{q-1}(|a|^{q}+|b|^{q})$
-by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality [29]. Suppose that
+by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality Rudin [29]. Suppose that
 $1>3\cdot 2^{2q}\chi^{q-1}\tau(\chi),$ (34)
@@
 So one can always choose a $\chi_{M}>0$ such that (39), (40) and (41) hold for $x\leq\chi_{M}$.
-Note that we also need the conditions $\log\delta^{-1}>1/\varepsilon$ and
-$s\geq(8+\varepsilon)\log\delta^{-1}\frac{8\log\delta^{-1}+1}{\varepsilon\log\delta^{-1}-1}$
+Note that we also need the conditions $\log\delta^{-1}>1/\epsilon$ and
+$s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log\delta^{-1}+1}{\epsilon\log\delta^{-1}-1}$
 to apply Lemma 1.
@@
 $\begin{split}&s_{i}(T)=\sum_{t=1}^{T}\mathbb{I}\left(A_{t}=i\right)\\ \leq&\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)+\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})\geq\theta_{1}\right).\end{split}$ (42)
-We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\varepsilon$ is equivalent to $t>\exp\left(\frac{1}{2\varepsilon}\right)$. Write
-$l_{0}:=\max\left\{\left\lfloor\exp\left(\frac{1}{2\varepsilon}\right)\right\rfloor,\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil\right\}.$ (30)
-Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\varepsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
+We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\epsilon$ is equivalent to $t>\exp\left(\frac{1}{2\epsilon}\right)$. Write
+$l_{0}:=\max\left\{\left\lfloor\exp\left(\frac{1}{2\epsilon}\right)\right\rfloor,\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\right\}.$ (43)
+Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\epsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
 $\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)$
-$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right)$
-$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\varepsilon)\log t^{2}\frac{8\log t^{2}+1}{\varepsilon\log t^{2}-1}\right).$
+$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right)$
+$\leq$ $l_{0}+\sum_{t=l_{0}+1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\text{ and }s_{1}(t-1)\geq(8+\epsilon)\log t^{2}\frac{8\log t^{2}+1}{\epsilon\log t^{2}-1}\right).$
 By Lemma 2, we have
@@
 Now we bound the second part of the right hand side of (LABEL:s_i_divide). Define
-$l_{1}:=\max\left\{l_{0},\left\lceil\frac{8+\varepsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}.$ (31)
+$l_{1}:=\max\left\{l_{0},\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}.$ (44)
 For sufficiently large $T$, we have
@@
 $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}.$
-For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
-$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\varepsilon)\log T^{2}\frac{8\log T^{2}+1}{\varepsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\varepsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (33)
-A Monte Carlo simulation example based the Pareto distribution of type III (see Sec.
-1.1) is shown in Fig.
-1.
-Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e., $\chi$ defined in (9) becomes sufficiently small such that, in the $\sup$ expression of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is nonzero and the growth rate of the right-hand side with $\theta$ becomes small).
-Figure 1: The performance of extended robust UCB
-2.2 Further Analysis and Comparison of the Extended Robust UCB
-Now we consider scenarios where specific values of $p$ and $q$ are chosen to further improve the efficiency of the extended robust UCB policy.
-Case 1.
-Suppose that the $p^{\prime}$-th order of moments of reward distributions exist with $p^{\prime}>2$ for all arms.
-Then we can choose $p$, $q$ such that $p=2q\leq p^{\prime}$.
-The inequality (16) can be refined to
-$\begin{split}\mathbb{E}[|Y_{i}-v_{q}|^{p/q}]&=\mathbb{E}Y_{i}^{2}+v_{q}^{2}-2\mathbb{E}v_{q}Y_{i}=\mathbb{E}Y_{i}^{2}-v_{q}^{2}=v_{p}-v_{q}^{2}\\ &\leq(C_{p,q}-1)v_{q}^{p/q}.\end{split}$ (34)
-Hence, we can refine $C^{\prime}$ from (13) to
-$C^{\prime}=\sqrt{12(C_{p,q}-1)},$ (35)
-which can be much smaller than the original definition (13).
-Let’s come back to the definition of $\chi_{M}$, which guarantees the three inequalities (26), (27) and (28) to hold.
-The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e., the constant $\chi_{M}$ will be larger and the regret upper bound will be smaller, which further implies that the policy may converge to the best arm faster.
-Case 2.
-In addition to the assumption in Case 1, we add an assumption that the $4$-th order moments of reward distributions exist as in Lattimore [21].
-Thus we can choose $p=4$ and $q=2$.
-In this case, our UCB in Definition 2 can be rewritten as
+For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
+$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (46)
+$\hfill\qed$
+Appendix F Proof of Near-Optimal Regret without Knowing $C_{p,q}$
+Recall all three required assumptions which are related to $C_{p,q}$:
+$1>$ $C^{\prime}x^{1-\frac{q}{p}},$ (47)
+$1>$ $3\cdot 2^{2q}x^{q-1}\tau(x),$ (48)
+$\Delta_{i}\geq$ $v_{q}^{1/q}x^{(q-1)/q}(B(x)^{1/q}+12^{1/q}).$ (49)
+From the previous deduction, we conclude that $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}$, where $l_{0}=\max\{\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}=O(\log T)$ and $l_{1}=\max\{l_{0},\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\rceil\}$.
+If we consider the case when replacing $C_{p,q}$ by $f(T)$ as long as T is large enough, $\chi_{M}$(an upper bound of $x$) here actually becomes a $T$-relevant value satisfying three inequalities, while the variation does not rely on the index of a certain arm since those requirements are established for all arms to attain.
+Now we only need to figure out the order of $\frac{1}{\chi_{M}}$ to prove our statement.
+For simplicity, we will discuss the three inequalities in turn. Recall that we use $f(T)$ instead of $C_{p,q}$ and focus on the case when $f(T)\geq C_{p,q}$, i.e. $T>t_{0}$ for some $t_{0}>0$.
 […]
-In this case $C_{p,q}$ degenerates to the upper bound on kurtosis. The function $B(x)$ defined in (23) becomes
-$B(x)=\frac{24\tau(x)}{1-48x\tau(x)}\left(1+C^{\prime}\sqrt{x}\right)+\frac{576\sqrt{x}\tau(x)}{1-48x\tau(x)},$
-and the three inequalities (26), (27) and (28) are refined to
-11 11 $>C^{\prime}\sqrt{x},$ (36)
-11 11 $>48x\tau(x),$ (37)
-$\Delta_{i}$ $\geq\sqrt{v_{2}x}\left(\sqrt{B(x)}+2\sqrt{3}\right).$ (38)
-Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (33) will be significantly decreased.
-Now we compare the extended robust UCB and the scale free algorithm proposed in Lattimore [21], under the same assumptions specified at the beginning of Case 2.
-We first address a flaw within Lemma 2 of Bubeck et al.
-[5].
-This lemma, crucial in Lattimore’s analysis [21] for deriving the regret upper bound of the scale free algorithm, is examined more closely below.
-For clarity and coherence with the notations employed in this paper, we present the following adapted version.
-A Flawed Statement [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$.
-Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d.
-random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $\mathbb{E}|X_{i}-\theta_{i}|^{p}=v_{p}$.
-Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$.
-Then with probability at least $1-\delta$,
-$\hat{\mu}\left(\{X_{i,j}\}_{j=1}^{s},k^{\prime}\right)\leq\theta_{i}+(12v_{p})^{1/p}\left(\frac{16\log({\rm e}^{1/8}\delta^{-1})}{s}\right)^{(p-1)/p}.$
-The last inequality in the proof of this statement in Bubeck et al.
-[5] assumes that $\exp(-k^{\prime}/8)\leq\delta$, and this is equivalent to $k^{\prime}\geq 8\log\delta^{-1}$.
-However, the definition of $k^{\prime}$ in this statement is
-$k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor.$
-To ensure that the inequality $\exp(-k^{\prime}/8)\leq\delta$ in their proof to hold, it is required that
-$\lfloor s/2\rfloor\geq 8\log\delta^{-1},$ (39)
-which is not assumed in the statement.
-As the statement is applied to the robust UCB, under the choice $\delta=t^{-2}$ and the definition $s:=s_{i}(t-1)$ as the number of times that arm $i$ was chosen before time $t$, Assumption (39) becomes
-$\lfloor s_{i}(t-1)/2\rfloor\geq 16\log t.$
-Hence the regret upper bound in Theorem 3 of Bubeck et al. [5] is at least $32\Delta_{i}\log T$.
-Below we give the corrected version of the above statement. Again, we will adopt notations introduced in this paper for consistency.
-Lemma 4.
-Let $\delta\in(0,1)$ and $p\in(1,2]$.
-Let $\{X_{i,j}\}_{j=1}^{s}$ be i.i.d.
-random variables with mean $\mathbb{E}X_{i}=\theta_{i}$ and (centered) $p$-th moment $\mathbb{E}|X_{i}-\theta_{i}|^{p}=v_{p}$, where $s$ is chosen such that $\lfloor s/2\rfloor\geq 8\log\delta^{-1}$.
-Let $k^{\prime}=\lfloor\min\{8\log({\rm e}^{1/8}\delta^{-1}),s/2\}\rfloor$.
-Then with probability at least $1-\delta$,
-$\hat{\mu}\left(\{X_{i,j}\}_{j=1}^{s},k^{\prime}\right)\leq\theta_{i}+(12v_{p})^{1/p}\left(\frac{16\log({\rm e}^{1/8}\delta^{-1})}{s}\right)^{(p-1)/p}.$
-Now we look into the results in Lattimore [21], whose Lemma 4 is a special version of the flawed Lemma 2 in Bubeck et al.
-[5].
-Lattimore’s Lemmas 5 and 6 in [21] both use the flawed version to estimate the regret upper bound as in the proof of their Theorem 2.
-We do not intend to give details about fixing all of them, but it is important to point out that, after applying the correct version (i.e., adding Assumption (39)) as in Lemma 4, the upper bound estimation of $\mathbb{E}[s_{i}(T)]$ in Lattimore [21] is at least (the constants introduced in their paper are explicitly calculated with notations translated here):
-$\begin{split}3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}\log({\rm e}^{1/8}\delta^{-1})+16\log\delta^{-1}.\end{split}$ (40)
-The next theorem shows that for properly chosen parameter $\varepsilon>0$, the upper bound given in (33) is also tighter than (40) over long-run if $\Delta_{i}$ is small, i.e., arms are hard to be distinguished (the low-discrimination case).
-Theorem 2.
-Choose any $8<\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$ ($\approx 44.158$) and let $\delta=t^{-2}$.
-The upper bound given in (40) of $\mathbb{E}[s_{i}(T)]$ is greater than that given in (33) for sufficiently large $T$ and sufficiently small $\Delta_{i}$.
-That is, by dividing both the upper bound estimations by $\log\delta^{-1}$ and letting $T\to+\infty$, we have
-$(8+\varepsilon)\left(\frac{8}{\varepsilon}+\max\left\{\frac{8}{\varepsilon},\frac{1}{\chi_{M}}\right\}\right)<3648\max\left\{\frac{(C^{\prime})^{2}}{12},\frac{v_{2}}{\Delta_{i}^{2}}\right\}+16$ (41)
-for sufficiently small $\Delta_{i}$.
-The regret upper bound of the scale free algorithm proposed in Lattimore [21] is thus larger than that of the extended robust UCB policy for $\varepsilon\in(8,\frac{280-16\sqrt{2}}{2\sqrt{2}+3})$.
-Proof.
-First note that, since $\Delta_{i}$ is sufficiently small and $\varepsilon$, $C^{\prime}$ are not related to $\Delta_{i}$, we only need to show (noticing that $\varepsilon>8$ implies $(8+\varepsilon)\frac{8}{\varepsilon}\leq 16$)
-$(8+\varepsilon)\frac{1}{\chi_{M}}<3648\frac{v_{2}}{\Delta_{i}^{2}}.$ (42)
+$\frac{1}{\chi_{M}}=O(f(T)^{\frac{q}{p-q}}).$ (50)
+[…]
+$\chi_{M}\textless[\frac{1}{c_{4}\cdot\tau(\chi_{M})}]^{\frac{q}{q-1}}.$ (51)
+After combining conditions given in (47) and (48), we see that (49) holds if
+$c_{0}\geq c_{1}\cdot\chi_{M}^{1-\frac{1}{q}}\left[c_{2}+c_{3}(\frac{\tau(\chi_{M})}{1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})})^{\frac{1}{q}}\right].$
+This indicates we only require $c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]^{\frac{1}{q}}\geq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})^{\frac{1}{q}}$ to achieve (49).
+Simple analysis shows that left side $\geq c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]$ while the right side $\leq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$.
+Thus we only require that
+$c_{0}\geq c_{0}c_{4}\chi_{M}^{q-1}\tau(\chi_{M})+c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$
+$=\chi_{M}^{1-\frac{1}{q}}[c_{0}c_{4}\chi_{M}^{q+\frac{1}{q}-2}\tau(\chi_{M})+c_{1}c_{2}+c_{1}c_{3}\tau(\chi_{M})].$
+By the fact that $\chi_{M}\leq 1$, the above inequality which is a stricter version of (49) holds if
+$c_{0}\geq\chi_{M}^{1-\frac{1}{q}}[(c_{0}c_{4}+c_{1}c_{3})\tau(\chi_{M})+c_{1}c_{2}],$
+i.e.
+$\chi_{M}\leq[\frac{c_{0}}{k\cdot\tau(\chi_{M})+b}]^{\frac{q}{q-1}},$ (52)
+where $k=c_{0}c_{4}+c_{1}c_{3},b=c_{1}c_{2}$.
+Furthermore, by comparing (51) with (52), they simultaneously hold if
+$\chi_{M}\leq\left[\frac{\min\{1,c_{0}\}}{\max\{k+b,c_{4}\}\cdot\tau(\chi_{M})}\right]^{\frac{q}{q-1}}=\left[d(1-(12(f(T)+1))^{\frac{q}{p}}\chi_{M}^{1-\frac{q}{p})}\right]^{\frac{q}{q-1}},$
+where $d=\frac{\min\{1,c_{0}\}}{\max\{k+b,c_{4}\}}$. Since $\chi_{M}<1$, we only require that
+$\chi_{M}\leq(\frac{d}{1+[12(f(T)+1)]^{\frac{q}{p}}d})^{\frac{1}{D}},~{}~{}~{}D=1-\max\{\frac{1}{q},\frac{q}{p}\}>0.$ (53)
+Finally, through comparison between (50) and (53), we have $\frac{1}{\chi_{M}}=O(f(T)^{\max\{\frac{q}{pD},\frac{q}{p-q}\}})=O(f(T)^{\frac{q}{pD}})$.
+By setting $l_{0}=\max\{t_{0},\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}$, we have $l_{1}=O(f(T)^{\frac{q}{pD}}\log T)$ and $\mathcal{R}_{\pi}^{\mathcal{F}}(T)=O(l_{0}+l_{1}+\frac{5\pi^{2}}{6})=O(f(T)^{\frac{q}{pD}}\log T)$.
+$\hfill\qed$
+Appendix G Proof of Theorem 2
+First note that, since $\Delta_{i}$ is sufficiently small and $\epsilon$, $C^{\prime}$ are not related to $\Delta_{i}$, we only need to show that (noticing that $\epsilon>8$ implies $(8+\epsilon)\frac{8}{\epsilon}\leq 16$)
+$(8+\epsilon)\frac{1}{\chi_{M}}<3648\frac{v_{2}}{\Delta_{i}^{2}}.$ (54)
 Equivalently, it is sufficient to show that
-$x^{\prime}:=\frac{(8+\varepsilon)\Delta_{i}^{2}}{3648v_{2}}<\chi_{M}$ (43)
+$x^{\prime}:=\frac{(8+\epsilon)\Delta_{i}^{2}}{3648v_{2}}<\chi_{M}$ (55)
 for sufficiently small $\Delta_{i}$.
 In order to prove the above, we need to show that (10), (11) and (12) are all true for $x=x^{\prime}$ based on the definition of $\chi_{M}$.
-For (36), let $x=x^{\prime}$, we have $1>\frac{C^{\prime}\Delta_{i}}{8}\sqrt{\frac{\varepsilon+8}{57v_{2}}}$, which is true for small $\Delta_{i}$.
+For (10), let $x=x^{\prime}$, we have $1>\frac{C^{\prime}\Delta_{i}}{8}\sqrt{\frac{\epsilon+8}{57v_{2}}}$, which is true for small $\Delta_{i}$.
 Furthermore, (11) becomes
-$1>6(8+\varepsilon)\Delta_{i}^{2}\frac{1}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(8+\varepsilon)v_{2}}},$
+$1>6(8+\epsilon)\Delta_{i}^{2}\frac{1}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(8+\epsilon)v_{2}}},$
 which also holds for sufficiently small $\Delta_{i}$. For (12), let $x=x^{\prime}$ and the inequality becomes
-$(\varepsilon+8)\left(\sqrt{6}\sqrt{\frac{v_{2}\left(\sqrt{57}(C^{\prime}+24)\Delta_{i}\sqrt{\frac{\varepsilon+8}{v_{2}}}+456\right)}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(\varepsilon+8)v_{2}}-6\Delta_{i}^{2}(\varepsilon+8)}}+\sqrt{3}\right)^{2}<912.$ (44)
-Taking the limit $\Delta_{i}\to 0$, (44) becomes $\left(6\sqrt{2}+9\right)(\varepsilon+8)<912$, which holds for $\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$.
-By continuity, we conclude that (43) holds for sufficiently small $\Delta_{i}$ if $\varepsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$.
-∎
-Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound by only a constant factor and a constant term as in the scale free algorithm of Lattimore [21] (see (40)).
-First, we restate the lower bound derived in Lattimore [21] as follows.
-Theorem 3.
-Let $H_{\kappa_{0}}$ be the set of distributions that has kurtosis less than $\kappa_{0}$.
-Assuming $\Delta>0$ and $\kappa_{0}\geq 7/2$.
-Suppose that $X\in H_{\kappa_{0}}$ has a mean of $\mu$, a positive variance of $\sigma^{2}$, and a kurtosis of $k$.
-Then
-$\begin{split}\inf\{{\rm KL}(X,X^{\prime}):&X^{\prime}\in H_{\kappa_{0}}\ \text{and}\ \mathbb{E}[X^{\prime}]>\mu+\Delta\}\\ \leq&\begin{cases}\min\{-\log(1-p),\frac{C_{1}\Delta^{2}}{\sigma^{2}}\},&\text{if}\ C_{0}\sqrt{k}(k+1)\frac{\Delta}{\sigma}<\kappa_{0}\\ -\log(1-p),&\text{otherwise}\end{cases},\end{split}$ (45)
-[…]
-Based on this, we can draw the following conclusion:
-Theorem 4.
-Let $\delta=t^{-2}$.
-Under case 2, the regret of the extended robust UCB policy applied on $H_{\kappa_{0}}$ ($\kappa_{0}\geq 7/2$) differs from the lower bound of the regret by a constant factor and a constant term.
-Proof.
+$(\epsilon+8)\left(\sqrt{6}\sqrt{\frac{v_{2}\left(\sqrt{57}(C^{\prime}+24)\Delta_{i}\sqrt{\frac{\epsilon+8}{v_{2}}}+456\right)}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(\epsilon+8)v_{2}}-6\Delta_{i}^{2}(\epsilon+8)}}+\sqrt{3}\right)^{2}<912.$ (56)
+Taking the limit $\Delta_{i}\to 0$, (56) becomes $\left(6\sqrt{2}+9\right)(\epsilon+8)<912$, which holds for $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$.
+By continuity, we conclude that (55) holds for sufficiently small $\Delta_{i}$ if $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$.
+$\hfill\qed$
+Appendix H Proof of Theorem 4
 If $\Delta_{i}$ of arm $i$ is large, that is, if (10) and (11) directly lead to (12), then $\mathbb{E}[s_{i}(T)]/\log T$ is upper bounded by a constant not related to $\Delta_{i}$.
@@
 where $C_{B}=B(1/{C^{\prime}}^{2})$ is a constant. Note that
-$\begin{split}\lim\sup\frac{\mathbb{E}[s_{i}(T)]}{\log T}&\leq 2(8+\varepsilon)\left(1+\max\left\{\frac{8}{\varepsilon},\frac{1}{\chi_{M}}\right\}\right)\\ &\leq 2(8+\varepsilon)\left(1+\max\left\{\frac{8}{\varepsilon},C_{B}\frac{v_{2}}{\Delta_{i}^{2}}\right\}\right).\end{split}$ (48)
-With the conclusion of (4) and Theorem 3, we finished the proof. ∎
-3 Light-Tailed Reward Disdtributions
-In this section, we formally and briefly discuss the past work of the first author of this paper for light-tailed reward distributions [24] and its comparisons with some subsequent work following this line.
-3.1 The UCB1-LT Policy
-The empirical mean is adopted in estimating the upper confidence bound for the class of light-tailed reward distributions.
-As before, let $\overline{X}_{i,s}$ denote the empirical mean $\frac{1}{s}\sum_{j=1}^{s}X_{i,j}$ for arm $i$, where $\{X_{i,j}\}_{j=1}^{s}$ forms the i.i.d.
-random reward sequence drawn from an unknown distribution $X_{i}\in\mathcal{F}$.
-Lemma 5.
-(Bernstein-type bound) For i.i.d.
-random variables $\{X_{i,j}\}_{j=1}^{s}$ drawn from a light-tailed distribution $X_{i}$ (with mean $\theta_{i}$) with a finite moment-generating function $M(u)$ over range $u\in[-u_{0},u_{0}]$ for some $u_{0}>0$.
-We have, $\forall\varepsilon>0$,
-$\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\geq\varepsilon)\leq\begin{cases}\exp(-\frac{s}{2\zeta}\varepsilon^{2}),&\varepsilon<\zeta u_{0}\\ \exp(-\frac{su_{0}}{2}\varepsilon),&\varepsilon\geq\zeta u_{0}\\ \end{cases},$ (49)
-where $\zeta>0$ satisfies $\zeta\geq\sup_{|u|\leq u_{0}}{M^{(2)}(u)}$. A similar bound holds for $\mathbb{P}(\overline{X}_{i,s}-\theta_{i}\leq-\varepsilon)$ by symmetry.
-The proof of the above lemma follows a similar argument as in Vershynin [35]. Using the Bernstein-type bound by Lemma 5, we propose the UCB1-LT policy as follows.
-1Input: $a_{1}\geq 8\zeta$, $a_{2}\geq a_{1}/(\zeta u_{0})$, $K$: the number of arms
-2 Initialize: $s_{i}\leftarrow 0$ and $\overline{X}_{i}\leftarrow 0$ for each arm $i$, $1\leq i\leq K$
-3 for $t\leftarrow 1$ to $T$ do
-4       for $i\leftarrow 1$ to $K$ do
-5             if $s_{i}=0$ then
-6                  Assign +$\infty$+ ∞ to the upper confidence bound
-7             end if
-8            else
-9                   if $\sqrt{\frac{a_{1}\log t}{s_{i}}}<\zeta u_{0}$ then
-10                         Compute the upper confidence bound as $\overline{X}_{i}+\sqrt{\frac{a_{1}\log t}{s_{i}}}$
-11                   end if
-12                  else
-13                         Compute the upper confidence bound as $\overline{X}_{i}+\frac{a_{2}\log t}{s_{i}}$
-14                   end if
-15
-16             end if
-17
-18       end for
-19      Choose arm $j$ that maximizes the upper confidence bound and obtain reward $x$
-20       $\overline{X}_{j}\leftarrow\frac{x+s_{j}\overline{X}_{j}}{s_{j}+1}$ and $s_{j}\leftarrow s_{j}+1$
-21 end for
-Algorithm 2 The UCB1-LT Policy
-The UCB1-LT policy considers two upper confidence bounds and alternatively uses one of them according to values of $t$ and $s_{i}(t-1)$.
-To minimize the theoretical regret upper bound, we can choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$.
-The following theorem shows that UCB1-LT achieves the logarithmic regret growth for light-tailed reward distributions.
-Theorem 5.
-For light-tailed reward distributions, the regret of the UCB1-LT policy satisfies the following inequality:
-$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{i:\theta_{i}<\theta_{1}}\Delta_{i}\left(\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}\log T+1+\frac{\pi^{2}}{3}\right),$ (50)
-Proof.
+$\begin{split}\lim\sup\frac{\mathbb{E}[s_{i}(T)]}{\log T}&\leq 2(8+\epsilon)\left(1+\max\left\{\frac{8}{\epsilon},\frac{1}{\chi_{M}}\right\}\right)\\ &\leq 2(8+\epsilon)\left(1+\max\left\{\frac{8}{\epsilon},C_{B}\frac{v_{2}}{\Delta_{i}^{2}}\right\}\right).\end{split}$ (59)
+With the lower bound in the non-parametric setting in Burnetas and Katehakis [6] and Theorem 3, the proof is finished immediately. $\hfill\qed$
+Appendix I Proof of Theorem 5
 Define
@@
 $\mathbb{E}[s_{i}(T)]\leq L_{0}+2\sum_{t=1}^{\infty}\sum_{s=1}^{t-1}\sum_{k=L_{0}}^{t-1}t^{-4}\leq\max\left\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\right\}+1+\frac{\pi^{2}}{3},$
-A direct substitution of the above into (1) completes the proof. ∎
-3.2 From UCB1-LT to $(\alpha,\psi)$-UCB
-After the establishment of UCB1-LT by Liu and Zhao [24], Bubeck and Cesa-Bianchi [4] subsequently proposed $(\alpha,\psi)$-UCB under a more general assumption on the moment generating functions.
-Specifically, the $(\alpha,\psi)$-UCB policy assumes that the distribution of reward $X_{i}\in\mathcal{F}$ satisfies the following condition: there exists a convex function $\psi$ defined on $\mathbb{R}$ such that for all $\lambda>0$,
-$\log\mathbb{E}\left[{\rm e}^{\lambda(X_{i}-\theta_{i})}\right]\leq\psi(\lambda)\text{ and }\log\mathbb{E}\left[{\rm e}^{\lambda(\theta_{i}-X_{i})}\right]\leq\psi(\lambda).$ (51)
-For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (51) becomes the well-known Hoeffding’s lemma [4].
-Since +$\infty$+ ∞ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT.
-By using the Legendre-Frenchel transform $\psi^{*}$ of $\psi$, defined as
-$\psi^{*}(\varepsilon):=\sup_{\lambda\in\mathbb{R}}\lambda\varepsilon-\psi(\lambda).$
-The upper confidence bound of $(\alpha,\psi)$-UCB at time $t$ is defined as
-$I_{i}(t):=\overline{X}_{i,s_{i}(t-1)}+(\psi^{*})^{-1}\left(\frac{\alpha\log t}{s_{i}(t-1)}\right).$ (52)
-Theorem 6 ([4]).
-Assume that the reward distributions satisfy (51). Then $(\alpha,\psi)$-UCB with $\alpha>2$ achieves
-$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{\theta_{i}<\theta_{1}}\left(\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\log T+\frac{\alpha}{\alpha-2}\right).$ (53)
-For the UCB1-LT policy, we can choose
-$\psi(\lambda)=\begin{cases}\frac{\zeta\lambda^{2}}{2},&\lambda\leq u_{0}\\ +\infty,&\lambda>u_{0}\end{cases}.$ (54)
-Then (51) becomes the same assumption as in UCB1-LT (see (6)). In this case, we have
-$\psi^{*}(\varepsilon)=\begin{cases}\frac{\varepsilon^{2}}{2\zeta},&|\varepsilon|<\zeta u_{0}\\ u_{0}(|\varepsilon|-\frac{\zeta u_{0}}{2}),&|\varepsilon|\geq\zeta u_{0}\end{cases}.$
-Thus the inverse $(\psi^{*})^{-1}(x)$ of $\psi^{*}$ for $x\geq 0$ will be
-$(\psi^{*})^{-1}(x)=\begin{cases}\sqrt{2\zeta x},&0\leq x<\frac{\zeta u_{0}^{2}}{2}\\ \frac{x}{u_{0}}+\frac{\zeta u_{0}}{2},&x\geq\frac{\zeta u_{0}^{2}}{2}\end{cases}.$
-Furthermore, if we choose $\alpha=4$, then the upper confidence bound defined in (52) becomes
-$I_{i}(t)=\overline{X}_{i,s_{i}(t-1)}+\begin{cases}\sqrt{\frac{8\zeta\log t}{s_{i}(t-1)}},&\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}\\ \frac{4\log t}{u_{0}s_{i}(t-1)}+\frac{\zeta u_{0}}{2},&\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}\end{cases}.$ (55)
-Notice that, if we choose $a_{1}=8\zeta$ and $a_{2}=\frac{8}{u_{0}}$ in UCB1-LT, we get the same index function under the condition $\frac{\log t}{s_{i}(t-1)}<\frac{\zeta u_{0}^{2}}{8}$.
-For the case $\frac{\log t}{s_{i}(t-1)}\geq\frac{\zeta u_{0}^{2}}{8}$, the upper confidence bound of $(\alpha,\psi)$-UCB is not larger than that of UCB1-LT and thus may achieve a smaller theoretical upper bound of regret as follows:
-$\mathcal{R}_{\pi}^{\mathcal{F}}(T)\leq\sum_{\theta_{i}<\theta_{1}}\left(r(\Delta_{i})\log T+2\right),$ (56)
-where
-$r(\Delta_{i})=\begin{cases}\frac{32\zeta}{\Delta_{i}},&\Delta_{i}<2\zeta u_{0}\\ \frac{8\Delta_{i}}{u_{0}(\Delta_{i}-\zeta u_{0})},&\Delta_{i}\geq 2\zeta u_{0}\end{cases}.$
-Note that this regret bound is better than (50) in Theorem 5 for UCB1-LT as $T$ becomes sufficiently large, as shown in the following lemma.
-Lemma 6.
-Choose the parameter $\alpha$ in $(\alpha,\psi)$-UCB such that $\alpha\leq 4$, and assume that Assumption (6) is satisfied and thus $\psi(\lambda)$ is chosen as in (54).
-Then the logarithmic regret upper bound in Theorem 6 has a leading constant no larger than that in Theorem 5, i.e.,
-$\sum_{\theta_{i}<\theta_{1}}\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\leq\sum_{\theta_{i}<\theta_{1}}\Delta_{i}\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}.$
-Proof.
+A direct substitution of the above into (1) completes the proof. $\hfill\qed$
+Appendix J Proof of Lemma 6
 We only need to show that
@@
 $x\max\left\{\frac{4a_{1}}{x^{2}},\frac{2a_{2}}{x}\right\}\geq\max\left\{\frac{32\zeta}{x},\frac{16}{u_{0}}\right\}=\begin{cases}\frac{32\zeta}{x},&0<x<2\zeta u_{0}\\ \frac{16}{u_{0}},&x\geq 2\zeta u_{0}\end{cases}.$
-Since $\frac{8\alpha\zeta}{x}\leq\frac{32\zeta}{x}$ as $\alpha\leq 4$ and $\frac{2\alpha x}{u_{0}(x-\zeta u_{0})}\leq\frac{16}{u_{0}}$ as $x\geq 2\zeta u_{0}$, (57) is proved. ∎
-Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments.
-Assume that the arms are offering rewards according to $19$ normal distributions.
-The means of reward distributions are $0.5,1.0,1.5,\dots,9.5$ with standard deviations all set to $60$.
-A Monte Carlo (MC) simulation with 100 runs and time period $T=50000$ is shown below.
-The horizontal axis in both figures is the time step $t$.
-The first figure only shows $T$ from $10000$ to $50000$ for better display, while the second one ranges from $1$ up to $50000$.
-The vertical axis of the left figure is the logarithmic regret averaged from the 100 MC runs (to approximate the expectation), while of the right one is the approximately expected time-average reward.
-The parameters for UCB1-LT and $(\alpha,\psi)$-UCB are respectively $u_{0}=1$ and $\alpha=2.5$ or $4$ with $\zeta=3600$ as their common parameter.
-These results reveal an interesting phenomenon: the actual performance of UCB-LT is similar to that of $(\alpha,\psi)$-UCB which has a better theoretical regret bound over long-run for $\alpha=4$.
-However, a smaller choice of $\alpha$ slightly improves the performance of $(\alpha,\psi)$-UCB over long-run as consistent with the theoretical bound in (53) whose leading constant of the logarithmic order decreases as $\alpha$ decreases.
-[Uncaptioned image][Uncaptioned image]
-Figure 2: Performance comparison between UCB1-LT and $(\alpha,\psi)$-UCB
-4 Conclusions
-In this paper, We have proposed two UCB policies, namely the extended robust UCB and UCB1-LT, dealing with the heavy-tailed and light-tailed reward distributions in the frequentist multi-armed bandit problems, respectively.
-The extended robust UCB policy aims at the class of heavy-tailed arm distributions and performs well in achieving the optimal logarithmic regret growth order.
-The policy takes advantage of a controlled relationship between different moments and discards the prerequisite knowledge of an upper bound of the moments.
-It is a generalization of the scale free algorithm proposed by Lattimore [21] with the assumption that $p=4$.
-In scenarios characterized by low discrimination and when $p=4$, our extended robust UCB policy exhibits a tighter regret upper bound in the long run.
-The achieved regret is merely a constant factor plus a constant term away from the established lower bound.
-The UCB1-LT is efficient for the class light-tailed reward distributions where the optimal logarithmic regret growth order is attained.
-Through numerical evaluations, we observe a comparable actual performance to its extension, $(\alpha,\psi)$-UCB.
-Since the upper confidence bound in UCB1-LT is easy to compute (with $O(1)$ time complexity), it provides a practical archetype of UCB policies for light-tailed reward distributions.
+Since $\frac{8\alpha\zeta}{x}\leq\frac{32\zeta}{x}$ as $\alpha\leq 4$ and $\frac{2\alpha x}{u_{0}(x-\zeta u_{0})}\leq\frac{16}{u_{0}}$ as $x\geq 2\zeta u_{0}$, (60) is proved. $\hfill\qed$

### v5

@@
 Abstract
-The multi-armed bandit (MAB) problem is a widely studied model in the field of operations research for sequential decision making and reinforcement learning.
-This paper mainly considers the classical MAB model with the heavy-tailed reward distributions.
-We introduce the extended robust UCB policy, which is an extension of the pioneering UCB policies proposed by Bubeck et al.
-[5] and Lattimore [22].
-The previous UCB policies require some strict conditions on the reward distributions, which can be hard to guarantee in practical scenarios.
-Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p>q>1$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of the UCB policies for the heavy-tailed reward distributions.
+The multi-armed bandit (MAB) problems are widely studied in fields of operations research, stochastic optimization, and reinforcement learning.
+In this paper, we consider the classical MAB model with heavy-tailed reward distributions and introduce the extended robust UCB policy, which is an extension of the results of Bubeck et al.
+[5] and Lattimore [22] that are further based on the pioneering idea of UCB policies [e.g.
+Auer et al.
+3].
+The previous UCB policies require some strict conditions on reward distributions, which can be difficult to guarantee in practical scenarios.
+Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p>q>1$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of UCB policies for heavy-tailed reward distributions.
 Furthermore, we achieve a near-optimal regret order without any knowledge of the reward distributions as long as their $p$-th moments exist for some $p>1$.
+Finally, we briefly present our earlier work on light-tailed reward distributions for a complete illustration of the amazing simplicity and power of UCB policies.
 1 Introduction
-The multi-armed bandit problem is an important topic in operations research where a player aims to identify and select the best arm to operate such that the long-term expected reward is maximized.
-The player’s action at each time involves balancing the trade-off between exploration and exploitation, i.e., choosing an arm to learn its statistics vs.
+As a prototype of reinforcement learning, a classical multi-armed bandit (MAB) model consists of a player and a bandit machine with $K$ arms.
+Under resource constraints, the player can only choose one arm to pull at each time step.
+The objective of the player is to identify and select the best arm so that the long-term expected reward is maximized.
+Clearly, the player’s action involves balancing the trade-off between exploration and exploitation, i.e., choosing an arm to learn its statistics vs.
 choosing an arm appearing to be best based on past observations.
-Under resource constraints, the way we acquire knowledge becomes crucial in maximizing the long-term reward rate in modern AI systems [31].
-Since knowledge can generally be treated as data, the struggle between exploration (sampling) and exploitation (consuming) as the prototypical dilemma in reinforcement learning is expressly reflected in an MAB model [28, 21, 3, 24, 5].
-Specifically, the original stochastic model of the multi-armed bandit problem considers a bandit machine with several arms; each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or agent).
+Specifically, each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or agent).
 Each reward distribution has a mean, say $\theta_{i}$.
 The player who has limited knowledge about the underlying reward distributions (no knowledge about the means $\{\theta_{i}\}$), has to choose only one of these arms to play and accrue a reward at each time step, given the reward history so far observed.
-The aim of the player is to maximize the reward rate that accrues as the time horizon $T$ grows.
+The player aims to maximize the reward rate that accrues as the time horizon $T$ grows.
 In many applications of online decision-making problems, such as Internet advertising and recommendation systems, the MAB model is widely assumed, with numerous algorithms proposed for efficient reinforcement learning, e.g.
 Zhao [38], Lattimore and Szepesvári [23].
-To address the key issue in the conflict between exploration and exploitation, we note that it is impossible for the player to exactly know the values of reward means within any finite horizon of time.
-The only information the player possesses is the (local) reward history, and the player has to learn from this to make a decision at each time step.
-Based on the limited historical information, the player has to choose between two options.
-The first option is to explore a suboptimal arm with a lower average reward, as there is always a possibility that the best arm may not give the best average reward in the past sample path.
-The second option is to exploit the arm with the highest average reward offered so far.
-Robbins [28] considered a simple two-armed bandit problem and showed that there is a policy to achieve the best average reward asymptotically as time goes to infinity.
-Many years later,Lai and Robbins [21] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the benchmark scenario in which the player already knows which arm has the highest reward mean from the very beginning.
-Given a set of $K$ arms with reward distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$ with means $\{\theta_{i}\}_{i=1}^{K}$ and a policy $\pi$, the regret ~$\tilde{\mathcal{R}}_{\pi}^{\mathcal{F}}(T)$ is defined as (without loss of generality, we can assume that $\theta_{1}\geq\theta_{i}$ for $1\leq i\leq K$)
+We point out that the way knowledge is acquired becomes crucial to maximizing the long-term reward rate in modern AI systems [31].
+Since knowledge can be treated as data in general, the struggle between exploration (sampling) and exploitation (consuming) as a fundamental dilemma is expressly reflected in the MAB model [28, 21, 3, 24, 5].
+To address the key issue in the conflict between exploration and exploitation, we emphasize on the fact that it is impossible for the player to know exactly the specific values of reward means within any finite horizon of time.
+Instead, the only information the player possesses is the (local) reward history, and the player has to learn from this to make a decision at each time step.
+Based on the limited historical information, the player has to choose between two specific options.
+The first option is to explore an apparently suboptimal arm with a lower average reward, as there is always a possibility that the truly best arm may not give the best average reward in the past sample path.
+The second option is to exploit an arm with the highest average reward offered so far.
+Robbins [28] considered a simple two-armed bandit model and showed that there is a policy to achieve the best average reward asymptotically as time goes to infinity.
+Many years later,Lai and Robbins [21] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the ideal scenario in which the player knows which arm has the highest reward mean from the very beginning.
+Given a set of $K$ arms with reward distributions $\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$ with means $\{\theta_{i}\}_{i=1}^{K}$ and a policy $\pi$, the regret $\mathcal{R}_{\pi}^{\mathcal{F}}(T)$ is defined as (without loss of generality, we can assume that $\theta_{1}\geq\theta_{i}$ for $1\leq i\leq K$)
 $\mathcal{R}_{\pi}^{\mathcal{F}}(T)=\sum_{\theta_{i}<\theta_{1}}\mathbb{E}[s_{i}(T)]\Delta_{i},$ (1)
@@
 $\Delta_{i}=\theta_{1}-\theta_{i},$ (2)
-and $s_{i}(T)$ denotes the number of times that arm $i$ was chosen during time steps 1 to $T$ under policy $\pi$.
-Certainly, a lower regret growth rate indicates better performance of a policy, and also a higher efficiency of learning.
-Furthermore, any sub-linear regret order implies the best expected time-average reward as $T\rightarrow\infty$.
-Strikingly, Lai and Robbins [21] proved an asymptotic $O(\log T)$ lower bound on the regret growth rate, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown.
-They also proposed an asymptotically optimal policy for a family of distributions, under which the best regret can be achieved asymptotically, both the logarithmic order and the leading coefficient.
+and $s_{i}(T)$ denotes the number of times that arm $i$ was chosen during time steps 1 to $T$ under the policy $\pi$.
+Certainly, a lower regret growth rate indicates better performance of a policy and also a higher efficiency of learning.
+Furthermore, any sublinear regret order with $T$ implies the asymptotic achievement of the maximum time-average reward as $T\rightarrow\infty$.
+Strikingly, Lai and Robbins [21] proved an asymptotic lower bound $O(\log T)$ on the regret growth rate, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown.
+They also proposed an asymptotically optimal policy for a family of distributions, under which the optimal regret can be achieved asymptotically, both the logarithmic order and the leading coefficient.
 Auer et al.
-[3] proposed a class of policies featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $[a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player.
+[3] proposed a class of policies featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player.
 Their UCB1 and UCB1-Tuned polices perform very well for the classical MAB problem with a theoretically proven regret bound for any finite time horizon.
-UCB policies are widely applied in machine learning development (e.g.
-recommendation systems [32, 17]) and various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [20], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [16].
-1.1 Heavy-Tailed MAB Model
-The class of heavy-tailed distributions is a concept in contrast to the light-tailed case.
-Light-tailed (or sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation, while heavy-tailed distributions have a relatively high probability of producing a large deviation.
+UCB policies are widely applied in machine learning developments (e.g.
+recommendation systems [32, 17]) and in various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [20], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [16].
+1.1 Light-Tailed and Heavy-Tailed MAB Models
+In the literature, MAB was first addressed under the assumption that all reward distributions are light-tailed, i.e., the sample mean converges to the true mean faster so learning becomes easier compared to the heavy-tailed case.
+Specifically, light-tailed (or sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation; while heavy-tailed distributions have a relatively high probability of producing a large deviation.
 There are several popular light-tailed probability distributions, such as Bernoulli, Gaussian, Laplacian, and Exponential.
-Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions Chareka [10].
-Formally, a random variable $X$ is called light-tailed if there exists some $u_{0}>0$ such that its moment-generating function is well-defined on $[-u_{0},u_{0}]$ [10]:
+Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions [10].
+Formally, a random variable $X$ is called light-tailed if there exists some $u_{0}>0$ such that its moment-generating function is well defined on $u_{0},u_{0}] [$10]:
 $M(u):=\mathbb{E}[\exp(uX)]<\infty,\ \ \forall|u|\leq u_{0}.$ (3)
@@
 Note that the above upper bound on $M(u)$ has a form of the moment-generating function of the Gaussian distribution, thus the name “sub-Gaussian” has been adopted.
-The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution can produce large values with high probabilities, leading to the nonexistence of its moment-generating function (even locally).
+The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution can produce large values with high probabilities, leading to nonexistence of its moment-generating function (even locally).
 Compared to the light-tailed class, the heavy-tailed one is harder to learn in terms of the rank of their means.
@@
 First, we assume that the player knows a constant $C_{p,q}$ for some $p$ and $q$ ($p>q>1$) such that [^1] By Jensen’s inequality, the existence of the $p$-th moment implies the existence of the $q$-th moment for any $1<q<p$.
-$\mathbb{E}[|X-\mathbb{E}X|^{p}]\leq C_{p,q}\left(\mathbb{E}[|X-\mathbb{E}X|^{q}]\right)^{p/q}.$ (5)
-The requirement of knowledge of $C_{p,q}$ can be removed with an arbitrarily small sacrifice of the regret order as seen in Sec.
+$\mathbb{E}[|X-\mathbb{E}X|^{p}]\leq C_{p,q}\left[\mathbb{E}[|X-\mathbb{E}X|^{q}]\right]^{p/q}.$ (5)
+We point out that the knowledge of specific values of $p$ and $q$ is not required.
+Furthermore, the requirement of knowledge of $C_{p,q}$ can also be removed with an arbitrarily small sacrifice of the regret order as seen in Sec.
 2.1.
 The above inequality is a direct generalization of the assumption of kurtosis in Lattimore [22] where $p=4$ and $q=2$.
-For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function Arnold [2]
+For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function [2]
 $F(x)=\begin{cases}1-\left(1+\left(\frac{x-a}{\sigma}\right)^{1/\gamma}\right)^{-1},&x>a\\ 0,&x\leq a\end{cases},$
-where $10<\gamma<$ 1 and $\sigma>0$. Simple computations show that
-$\mathbb{E}[|X-\mathbb{E}X|^{p}]=C(p,q,\gamma)\left(\mathbb{E}[|X-\mathbb{E}X|^{q}]\right)^{p/q}$
+where $0<\gamma<1$ and $\sigma>0$. Simple computations show that
+$\mathbb{E}[|X-\mathbb{E}X|^{p}]=C(p,q,\gamma)\left[\mathbb{E}[|X-\mathbb{E}X|^{q}]\right]^{p/q}$
 for $1<q<p<1/\gamma$, where
@@
 $C(p,q,\gamma)$ $=I(p,\gamma)/I(q,\gamma)^{p/q}.$
-However, the moment of this distribution is a function of $\sigma$.
-When $\sigma$ increases, the moment of any fixed order also increases, if it exists.
-The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment Bubeck et al.
-[5].
+Note that the moment of this distribution is a function of $\sigma$.
+When $\sigma$ increases, the moment of any fixed order also increases if it exists.
+The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment [5].
 Therefore, for an arm with an unknown $\sigma$, we cannot apply the robust UCB policy.
@@
 It is thus better to eliminate the need for any prior knowledge regarding the specific values or bounds on the moments.
-This motivates Lattimore [22] to consider a scale free algorithm.
+This motivates Lattimore [22] to consider the scale free algorithm.
+However, Lattimore’s algorithm requires specific values of $p$ and $q$ ($p=2q=4$) and cannot deal with the very heavy-tailed case (for $p$ close to $1$).
+Such requirements significantly limited the application of the scale free algorithm.
+In contrast, our algorithm only requires $p>1$ while $p$ can be unknown or unspecified.
+Furthermore, our algorithm achieves near-optimal regret even without any specific knowledge on the reward distributions including $C_{p,q}$.
+These results significantly broaden the applicability of the scale free algorithm.
 After we generalize Lattimore’s work, we will give a correction to a minor flaw of a lemma in Bubeck et al.
@@
 1.2 Related Work
-The theorems of the frequentist MAB have been developed for a long period of time and various milestones have been made since Lai and Robbins [21].
+The theory of frequentist MAB has been developed for a long period of time, and various milestones have been made since Lai and Robbins [21].
 In contrast to the Bayesian model, the frequentist MAB does not assume a priori knowledge of the initial probability or state of the system but learns the core parameters solely through the past sampling history, e.g., computing an upper confidence bound as a function of the frequency that an arm is selected and the values of the observed samples.
-This way of decision-making in balancing exploration with exploitation contrasts with the Bayesian model in which the tradeoff is usually addressed by dynamic programming for the transitional characteristics of the system or the observation model [13].
+This way of decision-making in the balance of exploration with exploitation contrasts with the Bayesian model in which the tradeoff is usually addressed by dynamic programming for the transitional characteristics of the system or the observation model [13].
 After the UCB1 policy by Auer et al.
@@
 They introduced the KL-UCB algorithm, which notably tightened the regret bound.
-This improvement was achieved under the assumption of bounded rewards with a known bound.
-Furthermore, KL-UCB policy achieves the asymptotic lower bound of regret growth for this class of rewards given in Burnetas and Katehakis [6].
+This improvement was achieved under the assumption of bounded rewards with known bounds.
+Furthermore, KL-UCB policy achieves the asymptotic lower bound of regret growth rate for this class of rewards given in Burnetas and Katehakis [6].
 Kaufmann et al.
-[19] combined the Bayesian method with the UCB class and proposed Bayes-UCB for bounded rewards with a known bound, achieving a theoretical regret bound similar to that of KL-UCB.
+[19] combined the Bayesian method with the UCB class and proposed Bayes-UCB for bounded rewards with known bounds, achieving a theoretical regret bound similar to that of KL-UCB.
 Numerical experiments demonstrated the high efficiency of learning by Bayes-UCB for Gaussian reward distributions.
@@
 These policies were neither implemented nor proved to achieve the logarithmic regret growth in the case of unbounded reward distributions.
-In 2011, the first author of this paper proposed UCB1-LT with a proven bound on the regret growth of the logarithmic order for the class of light-tailed reward distributions Liu and Zhao [24], filling the gap for the case of unbounded light-tailed reward distributions.
+In 2011, the first author of this paper proposed UCB1-LT with a proven logarithmic upper bound on the regret growth rate for the class of all light-tailed reward distributions [24], filling the gap for the case of unbounded light-tailed reward distributions.
 Together with the extended robust UCB for the heavy-tailed class focused in this paper, we provide a relatively complete picture of UCB policies for general reward distributions.
-After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the reward moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
+We will first present our UCB policy for the heavy-tailed class as the main content of this paper, followed by a brief presentation of UCB1-LT.
+These results demonstrate the simplicity and strong performance of UCB for all reward distributions.
+The UCB for the heavy-tailed case is of course more complicated because the player needs to extract more information from past observation in contrast to the simple form (e.g.
+sample mean) of UCB as in the light-tailed case.
+After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the (light-tailed) reward moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
 For the case of heavy-tailed reward distributions, we have mentioned that Bubeck et al.
@@
 The robust UCB policy uses different mean estimators rather than the empirical one and remarkably achieves the logarithmic regret growth rate.
-Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator Catoni [8].
-The truncated mean estimator requires the knowledge of an upper bound of origin moments, which is not very effective in the case that the reward mean is far away from zero.
+Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator [8].
+The truncated mean estimator requires the knowledge of an upper bound on the origin moments, which is not very effective in the case that the reward mean is far away from zero.
 The median-of-means estimator considers the (central) moment and thus patches up the deficiency of the truncated mean estimator: the mean estimation will not be affected when the reward distributions are transformed by translations.
@@
 Recently, Chen et al.
-[11] showed a result of applying Catoni mean estimator to the case of $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators.
+[11] showed the result of applying the Catoni mean estimator to the case of $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators.
 Finer estimations when extending the Catoni mean estimator to the case of $p<2$ are interesting for future investigations.
-Several recent papers also considered the heavy-tailed MAB model and made significant progresses.
+Several recent papers also considered the heavy-tailed MAB model and made significant progress.
 Wei and Srivastava [36] proposed the robust MOSS policy with logarithmic regret growth.
@@
 [12] considered the case where no knowledge of the moments is available except for the existence of moments of order greater than $1$, but a relatively strong assumption of the distribution (a larger probability mass on the negative semi-axis) was needed to achieve logarithmic regret.
-There are also policies not following the idea of UCB.
+There are also policies without following the idea of UCB.
 The $\epsilon$-greedy policy was well studied, and many policies were proposed using its idea, e.g., the constant $\epsilon$-decreasing policy [34], GreedyMix [9], and $\epsilon_{n}$-greedy [3].
@@
 Other recent progresses on variants of MAB, such as adversarial bandits, contextual bandits, and linear bandits, can be found in [23, 14, 18].
+In the following table, we list the main results found in the literature for classical MAB with heavy-tailed reward distributions for comparisons with ours.
+Papers Results Logarithmic Regret Bound Assumption
+Bubeck et al. [5] $\sum_{i:\Delta_{i}>0}\left(\frac{v}{\Delta_{i}}\right)^{1/\epsilon}\log T$ $\mathbb{E}[|X_{i}-\theta_{i}|^{1+\epsilon}]\leq v~\text{(known)},~\forall i$
+Lattimore [22] $\sum_{i:\Delta_{i}>0}\left((\kappa-1)\Delta_{i}+\frac{\sigma_{i}^{2}}{\Delta_{i}}\right)\log T$ $\mathrm{Kurt}[X_{i}]\leq\kappa~\text{(known)},~\forall i,~\sigma_{i}:=\sqrt{\mathrm{Var}(X_{i})}$
+Wei and Srivastava [36] $\sum_{i:\Delta_{i}>0}\log\left(\frac{T\Delta_{i}^{\frac{1+\epsilon}{\epsilon}}}{K}\right)\frac{1}{\Delta_{i}^{1/\epsilon}}$ $\mathbb{E}[|X_{i}-\theta_{i}|^{1+\epsilon}]\leq v~\text{(known)},~\forall i$
+Agrawal et al. [1] $\sum_{i:\Delta_{i}>0}\frac{\log T}{KL_{\inf}(\mathcal{F}_{i},\theta_{1})}$ $\mathbb{E}[|X_{i}-\theta_{i}|^{1+\epsilon}]\leq v~\text{(known)},~\forall i$
+Genalti et al.
+[12] $\sum_{i:\Delta_{i}>0}\left(\left(\frac{v}{\Delta_{i}}\right)^{1/\epsilon}+\frac{\Delta_{i}}{\mathbb{P}(X_{i}\neq 0)}\right)\log T$ $\mathbb{E}[X_{1}\mathds{1}_{\{|X_{1}|>M\}}]\leq 0,~\forall M\geq 0$
+Ours
+$\sum_{i:\Delta_{i}>0}\Delta_{i}\Big{(}144(1+C_{p,q})(6v_{q}^{\frac{1}{q}}+2^{q})(\Delta_{i}+\frac{1}{\Delta_{i}})\Big{)}^{M(p,q)}\log T$,
+$where\quad M(p,q)=1/[(1-\frac{1}{q})(1-\frac{q}{p})]>1$
+$v_{p}/v_{q}^{p/q}\leq C_{p,q}~\text{(known)},~1<q<p,~\forall i,~v_{l}:=\mathbb{E}[|X_{i}-\theta_{i}|^{l}]$
+$^{\dagger}KL_{\inf}(\eta,x):=\inf\{KL(\eta,\mathcal{F}):\mathbb{E}[|X|^{1+\epsilon}]\leq v\quad and\quad\mathbb{E}_{X\sim\mathcal{F}}[X]\geq x\}$
 2 Heavy-Tailed Reward Distributions
 2.1 The Extended Robust UCB Policy
-Consider $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with heavy-tailed distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$.
+Consider $K$ arms offering random rewards $\{X_{1},X_{2},\dots,X_{K}\}$ with heavy-tailed distributions $\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$.
 In contrast with the light-tailed class, a heavy-tailed distribution may yield a very high (or low) reward realization with a high probability.
@@
 $\frac{1}{s}\sum_{j=1}^{s}X_{i,j}\leq\theta_{i}+\left(\frac{3\mathbb{E}|X_{i}-\theta_{i}|^{r}}{\delta s^{r-1}}\right)^{\frac{1}{r}},$
-where $1<r\leq 2$. This bound is too loose to obtain a UCB policy for achieving logarithmic regret growth. To remedy this issue, we adopt the median-of-means estimator proposed in  Bubeck et al. [5].
+where $1<r\leq 2$.
+However, this bound is too loose to obtain a UCB policy for achieving logarithmic regret growth.
+To remedy this issue, we adopt the median-of-means estimator proposed in Bubeck et al.
+[5].
 Definition 1.
@@
 As mentioned in (5), we first assume that for any arm distribution $X_{i}\in\mathcal{F}$ with mean $\theta_{i}$, there is a known moment control coefficient $C_{p,q}$ ($1<q<p$) such that
-$\mathbb{E}[|X_{i}-\theta_{i}|^{p}]\leq C_{p,q}\mathbb{E}[|X_{i}-\theta_{i}|^{q}]^{r},$ (6)
+$\mathbb{E}[|X_{i}-\theta_{i}|^{p}]\leq C_{p,q}[\mathbb{E}[|X_{i}-\theta_{i}|^{q}]]^{r},$ (6)
 where $r=p/q$.
@@
 Let $k=\lceil 8\log\delta^{-1}\rceil,\chi=\frac{(8+\epsilon)\log\delta^{-1}}{s}$.
-Given an arbitrary $\epsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to +$\infty$+ ∞ if divided by 00)
+Given an arbitrary $\epsilon>0$, a moment order $p>1$, another moment order $1<q<p$, and the moment control coefficient $C_{p,q}$, define the upper confidence bound of the extended robust UCB as (value of the fraction is set to $\infty$ if divided by 0)
 $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta):=$
@@
 $C^{\prime}=\left(12(C_{p,q}+1)\right)^{q/p}.$ (7)
-Note that this upper confidence bound always exists, since the left-hand side of the above inequality tends to $\infty$- ∞ as $\theta\to-\infty$, while the right-hand side tends to +$\infty$+ ∞.
+Note that this upper confidence bound always exists, since the left-hand side of the above inequality tends to -$\infty$ as $\theta\to$-$\infty$, while the right-hand side tends to $\infty$.
 Lemma 2.
@@
     if $s_{i}=0$ then
-                  Set arm $i$’s upper confidence bound to +$\infty$+ ∞
+        Set arm $i$’s upper confidence bound to $\infty$
       end if
@@
 Then for any $\epsilon>0$, the regret $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ of the extended robust UCB policy $\pi^{*}$ has the logarithmic order with respect to the time horizon $T$.
+Specifically, we have $\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)\leq\sum_{i:\Delta_{i}>0}C_{i}(T)\Delta_{i}$ with
+$C_{i}(T):=\Big{[}36(1+\frac{8}{\epsilon})+(16+2\epsilon)\Big{(}144(1+C_{p,q})(6v_{q}^{\frac{1}{q}}+2^{q})(\Delta_{i}+\frac{1}{\Delta_{i}})\Big{)}^{M(p,q)}\Big{]}\log T+11,$
+where $M(p,q):=1/((1-\frac{1}{q})(1-\frac{q}{p}))$ and $T\geq\frac{9}{2\epsilon}+\frac{3}{2}.$
+From the above, we observe that the upper bound on regret increases with $C_{p,q}$ but decreases with $p$.
+This matches the intuition that a tighter known bound on the moment ratio or a higher moment order adopted makes the learning process easier due to less variation assumed in the samples.
+Furthermore, for any fixed $q$, the upper bound increases with $v_{q}$.
+This is also easy to explain since a larger moment indicates a larger deviation from a sample to the true mean.
+Meanwhile, if $q$ increases and becomes close to $p$, then the denominator of $M(p,q)$ will approach 0 and the upper bound becomes worse.
+This is because a larger value of $p/q$ offers more global information on moments with different orders to facilitate the learning of the unknown distributions.
+Also, as $\Delta_{i}\rightarrow 0$ for some $i$, the upper bound goes to infinity (since $M(p,q)>1$) because it will be harder to distinguish the best arm from arm $i$.
+Finally, if the arm number $K$ is increased with other parameters including $\Delta_{i}$ relatively fixed within the original orders, the upper bound is clearly increased linearly with $K$ due to the definition of regret.
 To prove Theorem 1, we just need to show the following core lemma to bound the time $s_{i}$ spent on each suboptimal arm $i$ by the extended robust UCB policy.
@@
 The expected number of times $\mathbb{E}[s_{i}(T)]$ that a suboptimal arm is chosen by the extended robust UCB policy $\pi^{*}$ has an upper bound of the logarithmic order with $T$.
-A Monte Carlo simulation example based on the Pareto distribution of type III (see Sec.
-1.1) is shown in Fig.
-1.
-Note that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e.
-$\chi$ defined in (24) becomes sufficiently small such that, within the $\sup$ operator of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, both the denominator is not zero and the growth rate of the right-hand side with $\theta$ becomes small).
+Specifically, we have $\mathbb{E}[s_{i}(T)]\leq C_{i}(T)$ where $C_{i}(T)$ is defined above.
+Finally, we conduct Monte Carlo simulation examples based on Laplace distributions, Student-t distributions, and Pareto distributions of type III (see Sec.
+1.1).
+From Fig.
+14, we observe that the regret rate grows linearly at the beginning to prepare the ground for ranking until at least one arm has a finite UCB (i.e., $\chi$ defined in (24) becomes sufficiently small such that, within the $\sup$ operator of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}},t^{-2})$, not only the denominator is nonzero but also the growth rate of the right-hand side with $\theta$ becomes small).
+In Fig.
+1, we compare the actual regret for different values of $C_{p,q}$ and observe that the regret becomes larger for larger $C_{p,q}$.
+In Fig 2, the regret becomes smaller as $p$ becomes larger.
+Furthermore, the regret increases with $v_{q}$ and $K$ as shown in Fig.
+3 and Fig.
+4, respectively.
+All these observations match our theoretical analysis on the upper bound above.
+(a) Laplace with $C_{p,q}=1.68\times 10^{3}$
+(b) Student-t with $C_{p,q}=2.09\times 10^{3}$
+Figure 1: Comparison for different $C_{p,q}$
+(a) Pareto III with $p=12$
+(b) Pareto III with $p=15$
+Figure 2: Comparison for different $p$
+(a) Laplace with $v_{q}=24$
+(b) Laplace with $v_{q}=384$
+Figure 3: Comparison for different $v_{q}$
+(a) Student-t with $K=2$
+(b) Student-t with $K=4$
+Figure 4: Comparison for different $K$
 Note that if $C_{p,q}$ is unknown, then the player can simply choose any increasing function $f$ with $f(t)\rightarrow+\infty$ as $t\rightarrow+\infty$ to replace $C_{p,q}$ and achieve a regret arbitrarily close to logarithmic order without requiring any knowledge of the reward moments (see Appendix F for a proof).
-Figure 1: The performance of extended robust UCB
 2.2 Further Analysis and Comparison of The Extended Robust UCB
-Now we consider scenarios of specific chosen $p$ and $q$ to further improve the efficiency of the extended robust UCB policy.
+Now we consider the specific chosen values of $p$ and $q$ to further improve the efficiency of the extended robust UCB policy.
 Case 1.
@@
 Then we can choose $p$ and $q$ such that $p=2q\leq p^{\prime}$.
-The inequality (29) can be refined to
+Inequality (29) can be refined to
 $\begin{split}\mathbb{E}[|Y_{i}-v_{q}|^{p/q}]=\mathbb{E}Y_{i}^{2}+v_{q}^{2}-2v_{q}\mathbb{E}Y_{i}=\mathbb{E}Y_{i}^{2}-v_{q}^{2}=v_{p}-v_{q}^{2}\leq(C_{p,q}-1)v_{q}^{p/q}.\end{split}$ (8)
@@
 which can be much smaller than the original definition (7).
-Now we look at the definition of $\chi_{M}$ in Appendix E that ensures the three inequalities (39), (40) and (41) to hold.
+Next, we look at the definition of $\chi_{M}$ in Appendix E that guarantees the three inequalities (39), (40) and (41) to hold.
 The refinement of $C^{\prime}$ leads to a smaller right-hand side for all these inequalities, i.e.
@@
 and the three inequalities (39), (40) and (41) are refined to
-11 11 $>C^{\prime}\sqrt{x},$ (10)
-11 11 $>48x\tau(x),$ (11)
+$1$ $>C^{\prime}\sqrt{x},$ (10)
+$1$ $>48x\tau(x),$ (11)
 $\Delta_{i}$ $\geq\sqrt{v_{2}x}\left(\sqrt{B(x)}+2\sqrt{3}\right).$ (12)
-Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (46) is significantly decreased.
+Under these refined estimations for $\chi_{M}$, the leading coefficient of $\log T$ in (46) is significantly reduced.
 Now we compare the extended robust UCB and the scale free algorithm proposed in Lattimore [22], under the same assumption specified at the beginning of Case 2.
@@
 For clarity and coherence with the notation employed in this paper, we present the following adapted version.
-A Flawed Statement  Bubeck et al.
+A Flawed Statement in Bubeck et al.
 [5]: Let $\delta\in(0,1)$ and $p\in(1,2]$.
@@
 The regret upper bound of the scale free algorithm proposed in Lattimore [22] is thus asymptotically larger than that of the extended robust UCB policy for $\epsilon\in(8,\frac{280-16\sqrt{2}}{2\sqrt{2}+3})$.
-Last, we show that the regret of the extended robust UCB deviates from the theoretical lower bound (in the non-parametric setting) by only a constant factor and a constant term as in the scale free algorithm of  Lattimore [22].
+Furthermore, we show that the regret of the extended robust UCB deviates from the theoretical lower bound (in the non-parametric setting) by only a constant factor and a constant term as in the scale free algorithm of Lattimore [22].
 First, we restate the lower bound derived in Lattimore [22] as follows.
@@
 $\begin{split}\inf\{{\rm KL}(X,X^{\prime}):&X^{\prime}\in H_{\kappa_{0}}\ \text{and}\ \mathbb{E}[X^{\prime}]>\mu+\Delta\}\\ \leq&\begin{cases}\min\{-\log(1-p),\frac{C_{1}\Delta^{2}}{\sigma^{2}}\},&\text{if}\ C_{0}\sqrt{k}(k+1)\frac{\Delta}{\sigma}<\kappa_{0}\\ -\log(1-p),&\text{otherwise}\end{cases},\end{split}$ (14)
-[…]
+where $C_{0}$, $C_{1}>0$ are universal constants and $p=\min\{\Delta/\sigma,\ 1/\kappa_{0}\}$.
 Based on this, we can draw the following conclusion:
@@
 Under Case 2, the regret of the extended robust UCB policy applied to $H_{\kappa_{0}}$ ($\kappa_{0}\geq 7/2$) differs from the lower bound of regret by a constant factor and a constant term.
+Furthermore, we give another upper bound that shows a sublinear regret in terms of arm number $K$. Recall the upper bound on regret in Theorem 1. We have
+$\mathcal{R}_{\pi^{*}}^{\mathcal{F}}(T)$ $=\sum_{i:\Delta_{i}>0}\Delta_{i}\mathbb{E}[s_{i}(T)]=\sum_{i:\Delta_{i}>0}\Delta_{i}[\mathbb{E}[s_{i}(T)]]^{1/M(p,q)}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}$
+$\leq\sum_{i:\Delta_{i}>0}\Delta_{i}\Big{[}\Big{(}M_{1}+M_{2}(p,q)(\frac{\Delta_{i}^{2}+1}{\Delta_{i}})^{M(p,q)}\Big{)}2\log T\Big{]}^{1/M(p,q)}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}$
+$\leq\sum_{i:\Delta_{i}>0}M_{3}(p,q)(\Delta_{i}^{2}+1)(\log T)^{1/M(p,q)}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}$
+$\leq M_{3}(p,q)(\log T)^{1/M(p,q)}\Big{[}\sum_{i:1\geq\Delta_{i}>0}2[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}+$
+$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\sum_{i:\Delta_{i}>1}2\Delta_{i}^{2}[\mathbb{E}[s_{i}(T)]]^{1-1/M(p,q)}\Big{]}$
+$\leq 2M_{3}(p,q)(\log T)^{1/M(p,q)}\Big{[}K^{1/M(p,q)}T^{1-1/M(p,q)}+\sum_{i:\Delta_{i}>1}\Delta_{i}^{2}T^{1-1/M(p,q)}\Big{]}$
+$\leq M_{3}(p,q)(\log T)^{1/M(p,q)}T^{1-1/M(p,q)}\Big{[}K^{1/M(p,q)}+\sum_{i:\Delta_{i}>1}\Delta_{i}^{2}\Big{]}$
+where $M_{1}=19+144/\epsilon,~M(p,q)=1/((1-1/q)(1-q/p))>1,~T\geq\frac{9}{2\epsilon}+\frac{3}{2}$, and
+$M_{2}(p,q)$ $=(8+\epsilon)\Big{(}144(1+C_{p,q})(6v_{q}^{\frac{1}{q}}+2^{q})\Big{)}^{M(p,q)},$
+$M_{3}(p,q)$ $=\Big{(}M_{1}+2M_{2}(p,q)\Big{)}^{1/M(p,q)}.$
+It is desirable to remove the dependency of regret bound on $\Delta_{i}$ as well and this will be considered in the future work.
 3 Light-Tailed Reward Distributions
-We point out that our policy and results also work for the light-tailed distributions since our construction depends only on the existence of moments.
-However, the leading constant of the logarithmic regret order can be improved by our past work in Liu and Zhao [24], which we formally and briefly discuss here with a comparison with some subsequent work following Liu and Zhao [24].
+We point out that our policy and results also work for the light-tailed reward distributions since our construction depends only on the existence of moments.
+However, as mentioned in Sec.
+1.2, the UCB could be significantly simplified for the light-tailed class where learning reward means becomes much easier.
+Furthermore, the leading constant of the logarithmic regret upper bound can also be greatly reduced as shown in [24].
+Therefore, we now formally but briefly discuss the work in [24] that extended the UCB policy for distributions with known finite supports in [3] to all light-tailed distributions.
+We will also compare our work with some subsequent work following [24].
+Together with above results for the heavy-tailed class, we see a complete picture of the beauty and power of UCB as well as what information needs to be extracted in its specific design based on how the samples represent the true means.
 3.1 The UCB1-LT Policy
@@
     if $s_{i}=0$ then
-                  Assign +$\infty$+ ∞ to the upper confidence bound
+        Assign $\infty$ to the upper confidence bound
       end if
@@
 Specifically, the $(\alpha,\psi)$-UCB policy assumes that the distribution of reward $X_{i}\in\mathcal{F}$ satisfies the following condition: there exists a convex function $\psi$ defined on $\mathbb{R}$ such that for all $\lambda>0$,
-$\log\mathbb{E}\left[{\rm e}^{\lambda(X_{i}-\theta_{i})}\right]\leq\psi(\lambda)\text{ and }\log\mathbb{E}\left[{\rm e}^{\lambda(\theta_{i}-X_{i})}\right]\leq\psi(\lambda).$ (17)
+[…]
 For example, if $X_{i}$ is bounded in $[0,1]$ as assumed in UCB1, one can choose $\psi(\lambda)=\frac{\lambda^{2}}{8}$ and (17) becomes the well-known Hoeffding’s lemma [4].
-Since +$\infty$+ ∞ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT.
+Since $\infty$ is allowed in the range of $\psi$, this assumption is more general than that in UCB1-LT.
 The Legendre-Frenchel transform $\psi^{*}$ of $\psi$ is defined as
@@
 $\sum_{\theta_{i}<\theta_{1}}\frac{\alpha\Delta_{i}}{\psi^{*}(\Delta_{i}/2)}\leq\sum_{\theta_{i}<\theta_{1}}\Delta_{i}\max\left\{\frac{4a_{1}}{\Delta_{i}^{2}},\frac{2a_{2}}{\Delta_{i}}\right\}.$
-Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments.
+Now we compare the actual performance between $(\alpha,\psi)$-UCB and UCB1-LT through numerical experiments shown in Fig.
+5.
 Assume that the arms are offering rewards according to $19$ normal distributions.
@@
 Since $s\geq(8+\epsilon)\log\delta^{-1}\frac{8\log(\delta^{-1})+1}{\epsilon\log(\delta^{-1})-1}$, a direct computation yields $\xi\leq 1/4$.
-Then using Hoeffding’s inequality for the tail of a binomial distribution Hoeffding [15], we have
+Then using Hoeffding’s inequality for the tail of a binomial distribution [15], we have
 $\mathbb{P}\left(\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)>\theta_{i}+\eta\right)\leq\mathbb{P}\left(\sum_{l=0}^{k-1}\mathbb{I}(\hat{\mu}_{l}\geq\theta_{i}+\eta)\geq\frac{k}{2}\right)\leq\exp(-2k(1/2-\xi)^{2})\leq\delta.$
-Similarly, the other inequality also holds by symmetry. $\hfill\qed$
+[…]
 Appendix C Proof of Lemma 2
@@
 Assume that $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)<+\infty$, i.e.
-$01-C^{\prime}\chi^{(p-q)/p}>$ 0 .
+$1-C^{\prime}\chi^{(p-q)/p}>0.$
 Define
@@
 $\theta_{i}\leq\hat{\mu}(\{X_{i,j}\}_{j=1}^{s},k)+\left(\frac{12\hat{\mu}(\{|X_{i,j}-\theta_{i}|^{q}\}_{j=1}^{s},k)}{\max\left\{0,1-C^{\prime}\chi^{(p-q)/p}\right\}}\right)^{1/q}\chi^{(q-1)/q}.$
-The proof is thus completed by the definition of $\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s},\delta)$. $\hfill\qed$
+[…]
 Appendix D Proof of Corollary  1
-This corollary is a direct consequence of Lemma 1 and the inequality in (29). $\hfill\qed$
+[…]
 Appendix E Proof of Lemma 3
@@
 $|a+b|^{q}\leq 2^{q-1}(|a|^{q}+|b|^{q})$
-by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality Rudin [29]. Suppose that
+by the convexity of $|x|^{q}$ ($q>1$) and Jensen’s inequality [29]. Suppose that
 $1>3\cdot 2^{2q}\chi^{q-1}\tau(\chi),$ (34)
@@
 $\Delta_{i}\geq$ $v_{q}^{1/q}x^{(q-1)/q}(B(x)^{1/q}+12^{1/q}).$ (41)
-[…]
+Observe that both $\tau(x)$ andB(x)B(x) decrease as $x$ decreases and
 $\lim_{x\searrow 0}x^{(q-1)/q}(B(x)^{1/q}+12^{1/q})=0.$
@@
 $\begin{split}&s_{i}(T)=\sum_{t=1}^{T}\mathbb{I}\left(A_{t}=i\right)\\ \leq&\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)+\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{i,j}\}_{j=1}^{s_{i}(t-1)},t^{-2})\geq\theta_{1}\right).\end{split}$ (42)
-We will bound the first part of the right hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\epsilon$ is equivalent to $t>\exp\left(\frac{1}{2\epsilon}\right)$. Write
+We will bound the first part of the right-hand side of (LABEL:s_i_divide) by Lemma 2. Note that $\log\delta^{-1}>1/\epsilon$ is equivalent to $t>\exp\left(\frac{1}{2\epsilon}\right)$. Write
 $l_{0}:=\max\left\{\left\lfloor\exp\left(\frac{1}{2\epsilon}\right)\right\rfloor,\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\right\}.$ (43)
-Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\epsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We thus have
+Also note that for sufficiently large $T$, the function $\log x^{2}\frac{8\log x^{2}+1}{\epsilon\log x^{2}-1}$ is increasing for $x\in[l_{0}+1,+\infty)$. We then have
 $\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)$
@@
 $\mathbb{E}\left[\sum_{t=1}^{T}\mathbb{I}\left(\tilde{\mu}(\{X_{1,j}\}_{j=1}^{s_{1}(t-1)},t^{-2})\leq\theta_{1}\right)\right]\leq l_{0}+\sum_{t=1}^{\infty}\frac{2}{t^{2}}\leq l_{0}+\frac{\pi^{2}}{3}.$
-Now we bound the second part of the right hand side of (LABEL:s_i_divide). Define
+Now we bound the second part of the right-hand side of (LABEL:s_i_divide). Define
 $l_{1}:=\max\left\{l_{0},\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}.$ (44)
@@
 $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}.$
-For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil$. Thus the above inequality becomes
-$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil,\left\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (46)
-$\hfill\qed$
+For sufficiently large $T$, we can assume that $l_{0}=\left\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\right\rceil$. Therefore, the above inequality becomes
+$\begin{split}\mathbb{E}[s_{i}(T)]&\leq\left\lceil(8+\epsilon)\log(T^{2})\frac{8\log(T^{2})+1}{\epsilon\log(T^{2})-1}\right\rceil\\ &+\max\left\{\left\lceil(8+\epsilon)\log(T^{2})\frac{8\log(T^{2})+1}{\epsilon\log(T^{2})-1}\right\rceil,\left\lceil\frac{8+\epsilon}{\chi_{M}}\log(T^{2})\right\rceil\right\}+\frac{5\pi^{2}}{6}.\end{split}$ (46)
+[…]
 Appendix F Proof of Near-Optimal Regret without Knowing $C_{p,q}$
@@
 From the previous deduction, we conclude that $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}$, where $l_{0}=\max\{\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}=O(\log T)$ and $l_{1}=\max\{l_{0},\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\rceil\}$.
-If we consider the case when replacing $C_{p,q}$ by $f(T)$ as long as T is large enough, $\chi_{M}$(an upper bound of $x$) here actually becomes a $T$-relevant value satisfying three inequalities, while the variation does not rely on the index of a certain arm since those requirements are established for all arms to attain.
+If we consider the case of replacing $C_{p,q}$ by $f(T)$ as long as $T$ is large enough, $\chi_{M}$(an upper bound of $x$) here actually becomes a $T$-relevant value satisfying three inequalities, while the variation does not depend on the index of a specific arm since those requirements are established for all arms to satisfy.
 Now we only need to figure out the order of $\frac{1}{\chi_{M}}$ to prove our statement.
@@
 $c_{0}\geq c_{1}\cdot\chi_{M}^{1-\frac{1}{q}}\left[c_{2}+c_{3}(\frac{\tau(\chi_{M})}{1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})})^{\frac{1}{q}}\right].$
-This indicates we only require $c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]^{\frac{1}{q}}\geq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})^{\frac{1}{q}}$ to achieve (49).
-Simple analysis shows that left side $\geq c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]$ while the right side $\leq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$.
-Thus we only require that
+This indicates that we only require $c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]^{\frac{1}{q}}\geq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})^{\frac{1}{q}}$ to achieve (49).
+A simple analysis shows the left side $\geq c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]$ while the right side $\leq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$.
+Therefore, we only require that
 $c_{0}\geq c_{0}c_{4}\chi_{M}^{q-1}\tau(\chi_{M})+c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$
@@
 where $k=c_{0}c_{4}+c_{1}c_{3},b=c_{1}c_{2}$.
-Furthermore, by comparing (51) with (52), they simultaneously hold if
+Furthermore, by comparing (51) with (52), they will simultaneously hold if
 $\chi_{M}\leq\left[\frac{\min\{1,c_{0}\}}{\max\{k+b,c_{4}\}\cdot\tau(\chi_{M})}\right]^{\frac{q}{q-1}}=\left[d(1-(12(f(T)+1))^{\frac{q}{p}}\chi_{M}^{1-\frac{q}{p})}\right]^{\frac{q}{q-1}},$
@@
 $\chi_{M}\leq(\frac{d}{1+[12(f(T)+1)]^{\frac{q}{p}}d})^{\frac{1}{D}},~~~D=1-\max\{\frac{1}{q},\frac{q}{p}\}>0.$ (53)
-Finally, through comparison between (50) and (53), we have $\frac{1}{\chi_{M}}=O(f(T)^{\max\{\frac{q}{pD},\frac{q}{p-q}\}})=O(f(T)^{\frac{q}{pD}})$.
-By setting $l_{0}=\max\{t_{0},\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}$, we have $l_{1}=O(f(T)^{\frac{q}{pD}}\log T)$ and $\mathcal{R}_{\pi}^{\mathcal{F}}(T)=O(l_{0}+l_{1}+\frac{5\pi^{2}}{6})=O(f(T)^{\frac{q}{pD}}\log T)$.
-$\hfill\qed$
+[…]
 Appendix G Proof of Theorem 2
@@
 $(\epsilon+8)\left(\sqrt{6}\sqrt{\frac{v_{2}\left(\sqrt{57}(C^{\prime}+24)\Delta_{i}\sqrt{\frac{\epsilon+8}{v_{2}}}+456\right)}{456v_{2}-\sqrt{57}C^{\prime}\Delta_{i}\sqrt{(\epsilon+8)v_{2}}-6\Delta_{i}^{2}(\epsilon+8)}}+\sqrt{3}\right)^{2}<912.$ (56)
-Taking the limit $\Delta_{i}\to 0$, (56) becomes $\left(6\sqrt{2}+9\right)(\epsilon+8)<912$, which holds for $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$.
-By continuity, we conclude that (55) holds for sufficiently small $\Delta_{i}$ if $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$.
-$\hfill\qed$
+[…]
 Appendix H Proof of Theorem 4
 If $\Delta_{i}$ of arm $i$ is large, that is, if (10) and (11) directly lead to (12), then $\mathbb{E}[s_{i}(T)]/\log T$ is upper bounded by a constant not related to $\Delta_{i}$.
-We are interested in the case where $\Delta_{i}$ is small, i.e., when (12) becomes the main restraint of $\chi_{M}$:
+We are interested in the case where $\Delta_{i}$ is small, i.e., when (12) becomes the main constraint of $\chi_{M}$:
 $\Delta_{i}=\sqrt{v_{2}\chi_{M}}\left(\sqrt{B(\chi_{M})}+2\sqrt{3}\right).$ (57)
@@
 $\begin{split}\lim\sup\frac{\mathbb{E}[s_{i}(T)]}{\log T}&\leq 2(8+\epsilon)\left(1+\max\left\{\frac{8}{\epsilon},\frac{1}{\chi_{M}}\right\}\right)\\ &\leq 2(8+\epsilon)\left(1+\max\left\{\frac{8}{\epsilon},C_{B}\frac{v_{2}}{\Delta_{i}^{2}}\right\}\right).\end{split}$ (59)
-With the lower bound in the non-parametric setting in Burnetas and Katehakis [6] and Theorem 3, the proof is finished immediately. $\hfill\qed$
+[…]
 Appendix I Proof of Theorem 5
@@
 $+\mathbb{P}\left(\frac{1}{s}\sum_{j=1}^{s}X_{1,j}\leq\theta_{1}-c(t,s)\right)+\mathbb{P}\left(\theta_{i}+2c(t,k)>\theta_{1}\right).$
-Choose $L_{0}=\lceil\max\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\}\rceil$, $\forall k\geq L_{0}$, we have
+Choose $L_{0}=\lceil\max\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\}\rceil$. We have, $\forall k\geq L_{0}$,
 $c(t,k)\leq$ $\max\left\{\sqrt{\frac{a_{1}\log t}{k}},\frac{a_{2}\log t}{k}\right\}\leq\max\left\{\sqrt{\frac{a_{1}\log t}{L_{0}}},\frac{a_{2}\log t}{L_{0}}\right\}$
@@
 $\mathbb{E}[s_{i}(T)]\leq L_{0}+2\sum_{t=1}^{\infty}\sum_{s=1}^{t-1}\sum_{k=L_{0}}^{t-1}t^{-4}\leq\max\left\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\right\}+1+\frac{\pi^{2}}{3},$
-A direct substitution of the above into (1) completes the proof. $\hfill\qed$
+[…]
 Appendix J Proof of Lemma 6
@@
 $x\max\left\{\frac{4a_{1}}{x^{2}},\frac{2a_{2}}{x}\right\}\geq\max\left\{\frac{32\zeta}{x},\frac{16}{u_{0}}\right\}=\begin{cases}\frac{32\zeta}{x},&0<x<2\zeta u_{0}\\ \frac{16}{u_{0}},&x\geq 2\zeta u_{0}\end{cases}.$
-Since $\frac{8\alpha\zeta}{x}\leq\frac{32\zeta}{x}$ as $\alpha\leq 4$ and $\frac{2\alpha x}{u_{0}(x-\zeta u_{0})}\leq\frac{16}{u_{0}}$ as $x\geq 2\zeta u_{0}$, (60) is proved. $\hfill\qed$
+[…]